### v1

# ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction

Abstract
Large language models (LLMs) are widely used as evaluators for open-ended tasks, while previous research has emphasized biases in LLM evaluations, the issue of non-transitivity in pairwise comparisons remains unresolved: non-transitive preferences for pairwise comparisons, where evaluators prefer A over B, B over C, but C over A. Our results suggest that low-quality training data may reduce the transitivity of preferences generated by the Evaluator LLM. To address this, We propose a graph-theoretic framework to analyze and mitigate this problem by modeling pairwise preferences as tournament graphs. We quantify non-transitivity and introduce directed graph structural entropy to measure the overall clarity of preferences. Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max), reflecting low overall clarity of preferences. To address this issue, we designed a filtering strategy, ELSPR, to eliminate preference data that induces non-transitivity, retaining only consistent and transitive preference data for model fine-tuning. Experiments demonstrate that models fine-tuned with filtered data reduce non-transitivity by 13.78% (from 64.28% to 50.50%), decrease structural entropy by 0.0879 (from 0.8113 to 0.7234), and align more closely with human evaluators (human agreement rate improves by 0.6% and Spearman correlation increases by 0.01).
ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction
[…]
1 Introduction
Figure 1: Non-Transitive Preferences in LLM-as-a-Judge for Pairwise Comparisons (e.g., $A\succ B$, $B\succ C$, $C\succ A$).
With the rapid advancement of large language model (LLM; OpenAI et al. 2024; Qwen Team 2025; Meta AI 2024a) technology, an increasing number of models have become available, making it essential to evaluate their capabilities for selecting the most suitable one. However, existing benchmarks such as MMLU Hendrycks et al. (2020) and HELM Liang et al. (2022) have been shown to be insufficient for capturing performance differences in open-ended tasks Zheng et al. (2024).
Given the lack of definitive answers in open-ended tasks, human expert evaluation is considered the gold standard. Yet, due to its high cost and limited scalability, the mainstream approach has shifted toward using LLM-as-a-Judge for efficient evaluation. Recent studies have demonstrated that powerful LLMs, such as GPT-4, can achieve high consistency with human judgments Zheng et al. (2024); Dubois et al. (2023). Among various approaches, pairwise comparison has emerged as the dominant paradigm, owing to its strong alignment with human preferences Samvelyan et al. (2024); Chen et al. (2024a); Li et al. (2023); Chiang et al. (2024); Liu et al. (2024); Liusie et al. (2024).
Despite these promising results, LLM-as-a-Judge still suffers from various biases—such as position, verbosity, conformity, and self-reinforcement bias—which compromise evaluation reliability Xu et al. (2024b); Zheng et al. (2024); Koo et al. (2024); Ye et al. (2025). A particularly underexplored yet critical issue is the non-transitivity of preferences generated by Evaluator LLMs—e.g., $A\succ B$, $B\succ C$, $C\succ A$, where $\succ$ denotes "is preferred to", i.e., $A\succ B$ means $A$ is preferred over $B$. An illustration is provided in Figure 1.
The first study to identify this issue Xu et al. (2025) analyzed inconsistencies in evaluation results when GPT-4 acted as the evaluator and performed pairwise comparisons using different baseline models. However, the study did not propose a method to mitigate this non-transitivity. Alarmingly, this issue is also inherited by specialized evaluators like JudgeLM, PandaLM, and Auto-J Zhu et al. (2025); Wang et al. (2024b); Li et al. (2024a), which are trained via knowledge distillation from foundation models. The distillation process may inadvertently propagate non-transitive judgment patterns to downstream evaluators.
We hypothesize that the presence of low-quality training data may impair the transitivity of the preferences generated by the Evaluator LLM. Many pairwise comparisons—particularly those from open-ended tasks—lack clear ground truth due to their subjective nature. Even human annotators often disagree, with inter-annotator agreement as low as 65.7% Li et al. (2023), indicating that such tasks inherently lack a universally accepted preference order. Consequently, training on ambiguous or low-quality comparisons can introduce or exacerbate non-transitive behavior in evaluators. This highlights the necessity of filtering unreliable data prior to model training.
In this work, we propose ELSPR, a simple yet effective method based on graph theory to assess and mitigate preference non-transitivity in Evaluator LLMs. We model multi-response pairwise comparisons as tournament graphs and filter out preference data that induce the non-transitivity of overall preferences. To further quantify the overall clarity of preference results, we introduce the use of two-dimensional structural entropy on directed graphs. Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max). Experimental results show that fine-tuning models on the filtered data leads to a 13.78% reduction in preference non-transitivity, a 0.0879 decrease in structural entropy, and improved alignment with human judgments—reflected in a 0.6% increase in human agreement rate and a 0.01 gain in Spearman correlation.
Specifically, our contributions are:
  • We introduce a graph-theoretic approach to systematically investigate the non-transitivity issue in preference data generated by Evaluator LLMs, revealing significant non-transitivity in the preference data produced by advanced LLMs.
  • We innovatively use the two-dimensional structural entropy of the directed graph to further evaluate the overall clarity of the preferences generated by large-scale LLMs.
  • We propose a effective filtering methodology that leverages tournament graphs to identify and remove preference data that induces non-transitivity, ensuring the retention of only consistent and transitive preferences. This enhances the reliability of Evaluator LLMs.
In addition, we released our code and datasets to facilitate further research into this issue [^1] https://github.com/yy0525/ELSPR.
2 Related Work
2.1 LLM-as-a-Judge and Its Non-Transitive Preferences
The prevailing LLM-as-a-Judge system predominantly operates through pairwise comparisons, as exemplified by frameworks such as VicunaEval, AlpacaEval, and Arena-HardChiang et al. (2023); Li et al. (2023, 2024b). Specifically, these systems collect responses generated by different LLMs to a given set of questions, and use advanced LLMs as judges to evaluate the preference order between two responses, assessing the relative strengths and weaknesses of different LLMs. Recent studies highlight that even advanced models like GPT-4 exhibit significant preference non-transitivity in LLM-as-a-Judge systems, which severely undermines the reliability of preference outcomes.
Despite prior efforts exploring six prompt templates to alleviate preference non-transitivity, improvements have been limited Xu et al. (2025). To address this critical issue, we investigate the underlying conditions that give rise to non-transitivity in LLM-as-a-Judge systems and propose a novel method to mitigate such inconsistencies.
Figure 2: ELSPR(Evaluator LLM training data Self-purification non-transitive Preferences via tournament graph Reconstruction) framework overview. The collected preference data are modeled as a directed graph. Strongly Connected Components (SCCs) are identified to analyze non-transitive structures, and each SCC is reconstructed into a DAG to filter out the specific preference data responsible for the overall non-transitivity.
2.2 Data Selection for LLM Fine-Tuning
Prior works Chen et al. (2024b); Ge et al. (2024); Li et al. (2024c) have highlighted the importance of filtering high-quality subsets from original datasets. Fine-tuning with such subsets not only enhances model performance but also reduces computational costs. However, this strategy remains underexplored in the context of the LLM-as-a-Judge task. Existing studies have primarily focused on scaling model size or improving prompt engineering, overlooking the potential benefits of training data quality. Our goal is to incorporate data filtering techniques into the LLM-as-a-Judge setting. We find that fine-tuning the LLM on preference data that has been filtered to induce non-transitivity can significantly reduce preference non-transitivity and improve overall preference clarity—even outperforming the teacher model.
3 Methodology
In this section, we present a graph-theoretic formulation of the LLM-as-a-Judge task using tournament graphs to analyze pairwise preferences from Evaluator LLMs. We then introduce quality analysis criteria for training data, including non-transitivity detection via Strongly Connected Components (SCCs) and preference clarity analysis via graph entropy. Finally, we propose a data filtering method to mitigate non-transitive preferences. Figure 2 outlines the overall framework, with implementation details provided in the following subsections.
3.1 Background
The typical LLM evaluation systems rely on the transitivity assumption, that is, if the evaluation results are $A\succ B$ and $B\succ C$, then it is assumed that $A\succ C$ However, evaluations in the real world often reveal non-transitivity preference cycles ($A\succ B\succ C\succ A$), which renders traditional ranking methods ineffective Xu et al. (2025). In order to overcome this fundamental limitation, we propose a tournament-based evaluation framework grounded in directed graphs, which facilitates the application of graph theory methods for a deeper analysis of the degree of non-transitivity in preferences and the overall clarity of the preferences.
Modeling preferences generated by LLM as a tournament graph. For each question $q_{i}\in Q$, given the response set $A_{i}=\{a_{1},a_{2},...,a_{n}\}$ from $n$ LLMs, we construct a tournament graph $G_{i}=(V_{i},E_{i})$ through the following procedure:
vertices: set $V_{i}=\{v_{1},v_{2},...,v_{n}\}$ corresponds to responses $A_{i}=\{a_{1},a_{2},...,a_{n}\}$.
Edges: Defined by preferences generated by LLM:
$E_{i}=\bigcup_{\begin{subarray}{c}1\leq j,k\leq n\\ j\neq k\end{subarray}}\begin{cases}v_{k}\rightarrow v_{j},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`win'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`lose'}\end{aligned}\\[10.00002pt] v_{j}\rightarrow v_{k},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`lose'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`win'}\end{aligned}\\[10.00002pt] v_{j}\leftrightarrow v_{k},&\text{otherwise}\end{cases}$ (1)
[…]
Considering the common position bias in Evaluator LLMs Wang et al. (2024a), we apply position swapping by comparing each response pair in both orders: $\mathcal{J}(a_{j},a_{k})$ and $\mathcal{J}(a_{k},a_{j})$. This ensures a more robust and balanced assessmentZheng et al. (2024). If the preferences differ across orders—indicating possible position bias—a bidirectional edge is added between the corresponding vertices to represent a ‘tie’.
3.2 Quality Analysis Criteria for LLM-as-a-Judge Training Data
Based on the directed graph representation of the tournament we constructed above (Section 3.1). We propose a quality analysis criteria for the training data of Evaluator LLM, which utilizes SCCs analysis to measure the degree of preference non-transitivity and employs two-dimensional structural entropy of directed graph structure entropy to evaluate the overall clarity of preferences.
3.2.1 Analysis of Non-Transitivity in Preferences Based on SCCs
The SCCs of a directed graph are maximal subgraphs in which, for any pair of vertices, there exists a directed path from one to the other and vice versa. This mutual reachability property aligns with the notion of preference cycles, which directly reflects non-transitivity in the preferences of the Evaluator LLM. A directed path from vertex $v_{i}$ to $v_{j}$ can be interpreted as a preference chain $a_{j}\succ\cdot\succ a_{i}$, while a reverse path from $v_{j}$ to $v_{i}$ corresponds to $a_{i}\succ\cdot\succ a_{j}$. The presence of both paths implies conflicting preferences and thus a violation of transitivity.
We use Tarjan algorithm to identify the SCCs of a directed graph. Tarjan algorithmTarjan (1972) is an efficient method for finding SCCs, based on depth-first search (DFS), with a time complexity of $O(N+M)$, where $N$ is the number of vertices and $M$ is the number of edges. Furthermore, since non-transitive relations occur between three or more vertices, we focus on components with a size greater than 2. Additionally, we filter out SCCs where there exists a bidirectional edge between every pair of vertices to eliminate cases where the relationship between vertices is a ‘tie’. For example, in the case of $(A=B),(B=C),(C=A)$, although $(A,B,C)$ forms an SCC, the preference relation is transitive. We denote the set of all SCCs in the directed graph $G$ as $SCCs(G)$,the following formula to obtain SCCs containing vertices with non-transitive relationships:
$\mathcal{S}_{\text{n-t}}=\left\{S\in SCCs(G)\,\middle|\,\begin{aligned} |S|>2\land\exists v_{j},v_{k}\in S,\\ (v_{j}\leftrightarrow v_{k})\notin E(S)\end{aligned}\right\}.$ (2)
To quantify the degree of non-transitivity in the preferences generated by the Evaluator LLM, we compute the ratio of non-transitive vertices to the total number of vertices across all directed graphs:
$\rho_{\text{non-trans}}=\frac{\sum_{q_{i}\in\mathbf{Q}}|S_{\text{n-t}}(G_{i})|}{\sum_{q_{i}\in\mathbf{Q}}|V_{i}|},$ (3)
where the numerator represents the total number of vertices belonging to non-transitive SCCs across all questions $\mathbf{Q}$, and the denominator represents the total number of vertices across all graphs. A higher value of $\rho_{\text{non-trans}}$ indicates a greater prevalence of non-transitive preferences.
3.2.2 Analysis of Preference Clarity Based the Entropy of Directed Graph Structures
The Shannon entropyShannon (1948) is a fundamental metric in information theory used to quantify the uncertainty associated with a random variable. It represents the expected level of uncertainty in a system: the greater the uncertainty, the higher the entropy. It is defined as:
$H(X)=-\sum p(x)\log p(x),$ (4)
where $p(x)$ denotes the probability distribution of the event $x$.
Structural entropy Li and Pan (2016) extends Shannon entropy to directed graphs, providing a measure of system uncertainty and the complexity of relationships within the graph. These concepts have been widely applied across various domains Zou et al. (2024); Duan et al. (2024); Peng et al. (2024); Hou et al. (2025). We introduce the Two-Dimensional Structural Entropy of a Directed Graph to analyze the overall clarity of preferences in Evaluator LLM.
Two-Dimensional Structural Entropy of a Directed Graph: The core idea of two-dimensional structural entropy is to partition the graph’s vertices into distinct communities and compute the complexity of both intra-community and inter-community relationships. This framework effectively quantifies the overall structural uncertainty and the complexity of relationships within the graph.
In this work, we innovatively adopt SCCs as the fundamental communities, motivated by properties from directed graph theory. Within an SCC, cyclic and non-transitive preference relations exist, while between SCCs, the preference relations are strictly transitive. This distinction allows us to quantitatively assess the overall preference clarity by computing the two-dimensional structural entropy of the tournament graphs constructed by Evaluator LLM.
It is important to note that SCCs composed of a single vertex inherently exhibit strict transitivity. The preference relations among such single-vertex components form a linear order that does not increase the complexity of the preference results. For example, if the original graph contains a preference $A\succ B$, the overall preference is clear. When new preferences $B\succ C$ and $A\succ C$ are added, the resulting linear order $A\succ B\succ C$ still remains clear and consistent — since the vertices $A$, $B$, and $C$ each form their own singleton SCC. This clarity reflects the inherent transitivity of singleton SCCs. Based on this observation, we exclude interactions among pure singleton SCCs when computing structural entropy. Instead, we retain only two types of interactions: (1) between singleton SCCs and multi-vertice SCCs, and (2) between multi-vertice SCCs. This selective inclusion ensures a more accurate and meaningful quantification of global structural complexity. Consequently, the closer the priority relations among vertices are to a linear order, the lower the resulting entropy.
The definition of the two-dimensional structural entropy of a directed graph based on SCC decomposition is as follows: Given a directed graph $G=(V,E)$, $n=|V|$ denote the number of vertices. Based on the SCC decomposition $SCCs(G)=\{SCC_{1},\dots,SCC_{L}\}$.
For a vertex $v\in V$, we use $d_{\text{in}}(v)$ and $d_{\text{out}}(v)$ to represent the in-degree and out-degree of $v$ in $G$, respectively. For any $SCC_{i}\subseteq SCCs$, define its volume as the total in-degree of its vertices:
$v(SCC_{i})=\sum_{v\in SCC_{i}}d_{\text{in}}(v).$ (5)
Accordingly, $v(G)$ denotes the total in-degree of the entire directed graph.
The two-dimensional structural entropy computation formula is as follows:
$H^{2}(G)=-\sum_{j=1}^{L}\frac{g_{j}}{v(G)}\log_{2}\frac{v(SCC_{j})}{v(G)}-\sum_{j=1}^{L}$ (6)
$\frac{v(SCC_{j})}{v(G)}\left(\sum_{v\in SCC_{j}}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\log_{2}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\right),$
where, $g_{j}$ denotes the total number of incoming edges to $V_{j}$ from external vertices. Specifically, it counts the edges between singleton and multi-vertice SCCs, as well as those among multi-vertice SCCs. It serves as a measure of the external influence on $V_{j}$.
Specifically, the normalized structural entropy for a graph $G$ is defined as:
$\tau(G)=\frac{H^{2}(G)}{\log n},$ (7)
where $H^{2}(G)$ represents the two-dimensional structural entropy of $G$, and $n$ denotes the number of vertices in the graph.
To evaluate the overall preference clarity of the Evaluator LLM across the set of questions $Q$, we compute the average normalized structural entropy, defined as follows:
$\tau_{\text{avg}}=\frac{\sum_{q_{i}\in\mathbf{Q}}\tau(G_{i})}{|\mathbf{Q}|},$ (8)
where $\tau(G_{i})$ is the normalized structural entropy for the graph $G_{i}$ corresponding to question $q_{i}$, and $|\mathbf{Q}|$ is the total number of questions in the set $\mathbf{Q}$. A higher value of $\tau_{\text{avg}}$ signifies that the preference relations generated by the Evaluator LLM are more complex, indicating a lower level of overall preference clarity.
3.3 Filtering Strategy for Preference Data That Induces Non-Transitivity
After decomposing a cyclic directed graph into multiple SCCs, if the cycles within each SCC can be eliminated and each SCC can be reconstructed as a DAG, then the entire directed graph can likewise be transformed into a Directed Acyclic Graph (DAG).
Building upon this observation, our framework transforms cyclic directed graphs into DAGs by converting each SCC into a DAG, preserving inter-SCC preference relations. This process ensures the global graph structure becomes acyclic without altering inter-component preferences. Subsequently, we derive transitive preference relations from the final DAG to filter Evaluator LLMs’ preference data, facilitating self-purification of the training data. The detailed procedure is as follows:
  1. 1.
    For each $SCC_{i}$ compute the in-degree $e_{i}^{\text{in}}$ for each vertex $v_{i}$, representing its global ‘win’ score. Vertices with higher in-degree scores are prioritized over those with lower scores. To reconstruct the internal edges within $SCC_{i}$, first remove all original edges between vertices within $SCC_{i}$. Then, for any pair $v_{i},v_{j}$: if $e_{i}^{\text{in}}>e_{j}^{\text{in}}$, add a directed edge $v_{j}\rightarrow v_{i}$; if $e_{i}^{\text{in}}=e_{j}^{\text{in}}$, add a bidirectional edge $v_{i}\leftrightarrow v_{j}$.
  2. 2.
[…]
  3. 3.
    The training data is traversed sequentially, and all data consistent with the correct preferences are added to the dataset "Cleaned".
See Appendix E for detailed algorithm information.
4 Experiment Setup
Figure 3: Comparison of data volumes between "Raw" and "Cleaned" training sets across datasets. The "Cleaned" dataset’s volume is approximately 80% of the "Raw" dataset for each dataset.
Model Helpful_Base Vicuna Oasst Koala Self-instruct
$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
Stronger LLMs
Qwen2.5-Max 63.68% 0.8047 75.36% 0.8448 64.29% 0.7883 71.52% 0.8296 64.97% 0.7803
Base models and Variants
Qwen-Base 82.83% 0.9221 78.93% 0.8912 83.36% 0.9138 80.96% 0.9097 81.46% 0.9123
Qwen-Raw 62.02% 0.7963 57.50% 0.8034 55.78% 0.7734 64.29% 0.8159 59.69% 0.7665
Qwen-Random 60.58% 0.7899 63.57% 0.8078 58.81% 0.7711 60.35% 0.8116 57.14% 0.7644
Qwen-Cleaned (ours) 44.85% 0.6997 43.93% 0.7262 46.96% 0.6942 48.53% 0.7149 48.98% 0.6802
Table 1: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw", "Random", and "Cleaned" training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base dataset.
Dataset: In this study, we conduct experimental validation using the AlpacaEval benchmark (Li et al., 2023). AlpacaEval is specifically designed to assess the overall capabilities of LLMs in open-domain, instruction-following tasks, covering a wide range of evaluation scenarios such as reasoning and text generation. The benchmark comprises five datasets: Helpful_Base, Oasst, Koala, Vicuna, and Self-Instruct.
Teacher Model Selection: We evaluated the performance of Qwen2.5-MaxQwen Team (2025) using the 2.5k human-annotated samples provided by AlpacaEval, employing the Chain-of-Thought Comparison (CoT Comparison) Li et al. (2023) . The detailed prompt template is provided in Appendix C.1. Qwen2.5-Max demonstrated exceptional performance in both human agreement rate, achieving an agreement level of 68.9%. Based on these results, we selected Qwen2.5-Max as our teacher model.
Experiment Details: We selected 21 representative LLMs from the AlpacaEval leaderboard and divided them into two groups based on the experimental design: 14 LLMs were used to collect preference data for constructing the training set, while the remaining 7 LLMs were employed to build the testing set for performance validation(see Appendix A for details). Throughout the experiment, we consistently applied the CoT Comparison prompt template for response generation, with the decoding temperature fixed at 0. Preference data for constructing the training set was generated by invoking the official Qwen2.5-Max API. The unfiltered Raw dataset was labeled as "Raw", and subsequently filtered according to the steps described in Section 3.3, resulting in the "Cleaned" dataset. The data distribution across datasets is illustrated in Figure 3. For comparative analysis, we fine-tuned models using both the "Raw" and "Cleaned" training sets derived from five datasets, ultimately obtaining a total of 10 fine-tuned models.
In the experiments, we used Qwen2.5-7B-Instruct as the base model and fine-tuned it using the Low-Rank Adaptation (LoRA) methodHu et al. (2022), with a rank of 8. The model was trained for 3 epochs, with a learning rate of $1\times 10^{-4}$ and a batch size of 16.
5 Results and Analysis
In this section, we validate the effectiveness of the proposed filtering method on five datasets. We utilize $\rho_{\text{non-trans}}$, as defined in Equation 3, to quantify the degree of preference non-transitivity in the evaluation of LLMs. A higher value of $\rho_{\text{non-trans}}$ indicates a greater extent of non-transitive preferences. Additionally, we employ $\tau_{\text{avg}}$, as defined in Equation 8, to assess the overall clarity of preferences, where a lower value reflects improved preference clarity. Furthermore, we provide a comprehensive measure of the alignment between model preferences and human preferences by calculating the human agreement rate and the spearman correlation coefficient between the model’s evaluations and human assessments. Additionally, we investigate two scenarios in which Evaluator LLMs are prone to exhibit non-transitive preferences, see Appendix D.
5.1 Main Results
5.1.1 Analysis of Non-Transitivity and Clarity in Preferences
As shown in Table 1, across all testing sets, models fine-tuned on the "Cleaned" training set exhibit the lowest degree of preference non-transitivity and the highest clarity of preferences, surpassing even the teacher model Qwen2.5-Max. This highlights the effectiveness of our proposed data filtering method. Additionally, models fine-tuned on the "Raw" dataset perform comparably to the teacher model Qwen2.5-Max, demonstrating that data distillation can indeed enhance a model’s evaluation capability relative to its original performance. However, due to the inherent preference non-transitivity of the teacher model, direct distillation may learn low-quality preference data, thereby limiting the ability of the distilled model to surpass the teacher model and resulting in similar performance.
To validate that models fine-tuned on the "Cleaned" training set achieve lower preference non-transitivity and higher preference clarity even on "unseen" questions, we performed cross-validation by testing each fine-tuned model on five testing sets. As shown in Table 2, models fine-tuned on Cleaned data show a 13.78% average reduction in non-transitivity and a 0.0879 decrease in normalized structural entropy compared to those trained on Raw data, indicating clearer overall preferences. Detailed results are in Appendix B.3. Notably, even advanced Evaluator LLMs (e.g., Qwen2.5-Max at 67.96%) exhibit high non-transitivity and entropy (0.8591 and 0.8095, respectively), reflecting limited preference clarity.
Training Set $\rho_{\text{non-trans}}$ $\tau_{\text{avg}}$
Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
Helpful_Base 65.95% 50.33% -15.62% 0.8204 0.7173 -0.1031
Vicuna 61.56% 50.96% -10.60% 0.8063 0.7367 -0.0696
Oasst 62.61% 48.73% -13.88% 0.7966 0.7101 -0.0865
Koala 64.05% 48.65% -15.40% 0.8044 0.7177 -0.0867
Self-instruct 67.23% 53.84% -13.39% 0.8286 0.7351 -0.0935
Average 64.28% 50.50% -13.78% 0.8113 0.7234 -0.0879
Table 2: Comparison of average preference non-transitivity and average clarity between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct across testing sets. The leftmost column indicates the source dataset used for training, while other columns show model average performance on five testing sets. For example, the "Raw" column under Helpful_Base shows results for unfiltered data, and the "Cleaned" column reflects filtered data.
5.1.2 Human Agreement Rate and Spearman Correlation Analysis.
We evaluate the alignment of fine-tuned Evaluator LLMs with 2.5k human-annotated preference labels from the AlpacaEval dataset, which contains around 650 questions with two candidate responses each, annotated by four human evaluators Li et al. (2023). The questions are sourced from Helpful_Base, Oasst, Koala, Vicuna, Self-Instruct, and Alpaca.
Human Agreement Rate: This metric quantifies the alignment between Evaluator LLMs and human annotators’ majority preferences. For each instruction, four predictions are generated by the Evaluator LLM. Using a leave-one-out approach, each prediction is compared against the majority vote of the remaining three human annotations, excluding the corresponding human annotation. The final score is computed as the average accuracy across all four comparisons per instruction.
Spearman Correlation:This metric evaluates the Spearman correlation between rankings based on Evaluate LLMs’ preferences and those derived from human preferences.
As shown in Table 3, models fine-tuned on the corresponding "Cleaned" training sets of Vicuna, Oasst, koala and Self-instruct exhibit higher human agreement rate compared to those fine-tuned on the "Raw" training sets, with improvements ranging from 0.6% to 1.2%. Additionally, models fine-tuned on the corresponding "Cleaned" training sets of Helpful_Base, Vicuna, and Self-instruct exhibit higher Spearman correlation compared to those fine-tuned on the "Raw" training sets. Notably, on the Helpful_Base "Cleaned" training set achieves the largest improvement, with the Spearman correlation increasing from 0.93 to 0.97.
5.2 Ablation Experiments
Ablation of Different Base Models: We conducted repeated experiments using the LLaMA3.1-8B-Instruct model as the base model. As shown in Appendix B.2, models trained on the "Cleaned" training set outperform those trained on the "Raw" training set and also surpass the advanced open-source Evaluator LLM Auto-J 13B. We further validated that this superiority holds true even for "unseen" questions, with detailed results provided in Appendix B.3, further highlighting the importance of high-quality data in enhancing model performance.
Ablation of Different Data Filtering Methods: As shown in Figure 3, the "Cleaned" training set constitutes approximately 80% of the "Raw" training set. To validate the effectiveness of our data filtering method, we randomly filtered 20% of the training data and fine-tuned the Qwen2.5-7B-Instruct model using the remaining 80%, denoted as Qwen-Random. As shown in Table 1, the overall performance is similar to the model trained using the "Raw" training set, and showed higher preference non-transitivity on the Vicuna testing set. This comparison underscores the effectiveness of our targeted filtering strategy in enhancing data quality.
Ablation of Different Prompt Forms: We conducted additional experiments using a prompt format that includes ‘tie’ preference outcomes. The detailed prompt template is provided in Appendix C.2, and the experimental setup and results are presented in Appendix B.1, further validating the effectiveness of our data filtering method. Notably, when using prompts that include ‘tie’ preference outcomes, Evaluate LLMs exhibit higher preference non-transitivity and lower overall preference clarity compared to prompts without ‘tie’ outcomes. This finding is consistent with the results reported in prior work Xu et al. (2025).
Training Set Human Agreement Rate Spearman Correlation
Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
Helpful_Base 66.9% 66.9% +0.0% 0.93 0.97 +0.04
Vicuna 65.2% 65.8% +0.6% 0.97 0.98 +0.01
Oasst 66.4% 67.6% +1.2% 0.93 0.93 +0.00
Koala 66.4% 66.9% +0.5% 0.98 0.98 +0.00
Self-instruct 67.8% 68.3% +0.5% 0.98 1.00 +0.02
Average 66.5% 67.1% +0.6% 0.96 0.97 +0.01
Table 3: Comparison of Human Agreement Rate and Spearman Correlation between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct. The structure matches Table 2.
6 Conclusion
In this paper, we present a detailed analysis of preference non-transitivity in pairwise evaluations by LLM-as-a-Judge using graph theory, while measuring overall clarity through the two-dimensional structural entropy of directed graphs. By filtering low-quality training data that induce non-transitive preferences, our method effectively reduces preference non-transitivity, enhances overall preference clarity, and improves alignment with human judgments. This principled framework contributes to more reliable automated evaluation with LLMs and offers valuable insights for future research on preference consistency and robust benchmarking.
Limitations and Future Work
Although we have validated the effectiveness of our filtering approach on five datasets from AlpacaEval, it still cannot fully cover all real-world scenarios, and further validation across a broader range of domains is needed. Additionally, while our method has successfully reduced the issue of preference intransitivity, it has not been completely eliminated. Further efforts to minimize preference intransitivity remain an important direction for future work. Lastly, our study primarily focused on pairwise comparison tasks in single-turn dialogues, exploring multi-turn dialogue tasks presents an intriguing avenue for future research.
Ethics Statement
This research focuses on analyzing and mitigating non-transitive preferences in LLMs when used as evaluators for open-ended tasks. The methodology involves processing publicly available data and conducting computational experiments using existing datasets and open-source frameworks, without involving human subjects or sensitive personal information. Therefore, no specific ethical approval was required for this study.
Appendix A LLM Details
This section provides a detailed overview of the models used to construct the training and testing sets in the experiment. We selected a total of 21 LLMs, covering advanced closed-source models, large-scale and small-scale open-source models, as well as dense and Mixture-of-Experts (MoE) architectures. Specifically, the selection includes the following:
  • OpenAI: OpenAI model series OpenAI et al. (2024).
  • Anthropic: Claude model seriesAnthropic (2023, 2024).
  • Mistral: Mistral model seriesJiang et al. (2024, 2023).
  • Alibaba: Qwen model seriesQwen Team (2024a, b).
  • Meta: LLaMA model series Meta AI (2024a, b); Touvron et al. (2023).
  • Yi: Yi model seriesAI et al. (2025).
  • Google: Gemini model seriesTeam et al. (2024).
  • Other Notable Models: Vicuna 33BChiang et al. (2023), Tulu 2+DPO 70BIvison et al. (2023), and WizardLM 13BXu et al. (2024a).
The pre-generated outputs for each model on the AlpacaEval dataset were sourced from the AlpacaEval leaderboard.
Appendix B Additional Experimental Results
B.1 Experimental Results of Different Prompt Forms.
We repeated the experiment using the CoT Comparison (Tie Allowed) prompt template. Figure 4 illustrates the distribution of dataset sizes before and after filtering across different training sets. The test results are presented in Table 5.
Across five testing sets, models fine-tuned on the "Cleaned" training set exhibit the highest preference clarity. In four of these testing sets (excluding Vicuna), these models also demonstrate the lowest preference non-transitivity, outperforming models fine-tuned on the "Raw" training set and the teacher model Qwen2.5-Max. We observe that models fine-tuned on the "Raw" training set, as well as the teacher model Qwen2.5-Max, exhibit a higher rate of ties in preferences on the Vicuna testing set. The prevalence of such ties—often arising from SCCs formed by cases like $(A=B)$, $(B=C)$, and $(C=A)$—significantly reduces the degree of preference non-transitivity when these SCCs are removed. As shown in Table 4, we compare the probability of ties among three models on the Vicuna testing set. Furthermore, the $\tau_{\text{avg}}^{\downarrow}$ values of models fine-tuned on the "Cleaned" training set indicate a higher overall clarity in preferences.
Figure 4: Comparison of Data Volumes Between "Raw" and "Cleaned" Training Sets Across Different Datasets (Using the CoT Comparison (Tie Allowed) Prompt Template)
Model tie probability
Qwen2.5-Max 7.59%
Qwen-Raw 10.06%
Qwen-Cleaned 6.13%
Table 4: Comparison of the probability of ties for models on the Vicuna testing set.
B.2 Experimental Results of Different Base Model.
We repeated the experiment using the LLaMA3.1-8B-Instruct model as the base model. The test results are presented in Table 6. We also compared our results with the advanced open-source Evaluator LLM Auto-J 13B, which is fine-tuned based on the LLaMA-2-13B-Chat model.
Due to the limited context length of 2048 tokens supported by the advanced open-source evaluators PandaLM and JudgeLM, which is insufficient to cover the evaluation tasks in our testing set, we did not include them in our experiments.
B.3 "Unseen" question validation
We perform a detailed comparison of the performance of models fine-tuned on different training sets and the advanced Evaluator LLM on various testing sets:
  • Results for models based on Qwen2.5-7B-Instruct are presented in Table 7.
  • Results for models based on LLaMA3.1-8B-Instruct are presented in Table 8.
Appendix C Evaluate Prompts
C.1 CoT Comparison
The CoT Comparison prompt template, identical to AlpacaEval 2.0 Li et al. (2023), is detailed in Table 9.
C.2 CoT Comparison (Tie Allowed)
The CoT Comparison (Tie Allowed) prompt template, as described in (Xu et al., 2025), is detailed in Table 10.
Appendix D Under What Scenarios Are Evaluator LLMs Prone to Non-Transitive Preferences?
Evaluator LLMs, much like human judges, tend to exhibit non-transitive preferences under certain conditions.
First, when the quality differences between responses fall below the Evaluator LLM’s just noticeable difference (JND) threshold (a concept widely used in psychophysics to describe the minimum difference perceptible to an observer Stern and Johnson (2010)) the model may "lose evaluative focus." Analogous to how humans struggle to differentiate between highly similar colors on a gradient from yellow to orange, the Evaluator LLM may perceive adjacent responses as "almost equally good." While individual comparisons appear consistent, cumulative evaluations can yield preference cycles (e.g., $A\succ B$, $B\succ C$, $C\succ A$), reflecting underlying non-transitivity.
Second, when responses differ significantly in stylistic dimensions, the Evaluator LLM may exhibit indecision due to the inherent conflict in multi-dimensional evaluation. Consider an open-ended question evaluated along three criteria: rigor, vividness, and richness:
  • Response A is the most rigorous, moderately vivid, and least rich.
  • Response B is the richest, moderately rigorous, and least vivid.
  • Response C is the most vivid, moderately rich, and least rigorous.
If overall preference is defined as outperforming another response in at least two out of three dimensions, instances such as $A\succ B$, $B\succ C$, and $C\succ A$ may arise—resembling the classic voting paradox, also known as the Condorcet paradox Fishburn (1974). This reflects a fundamental limitation in aggregating preferences across conflicting criteria: when each response possesses a distinct advantage, constructing a consistent global ranking becomes infeasible. The problem is especially pronounced in open-ended evaluations, where stylistic variance (e.g., concise and rigorous analysis vs. vivid and elaborate narratives) can lead to non-transitive outcomes in Evaluator LLM preferences.
Appendix E Algorithm Detail
Algorithm 1 Filtering Strategy for Preference Data That Induces Non-Transitivity
1:Cyclic directed graph $G=(V,E)$
2:Cleaned dataset with transitive preference relations
3:Decompose $G$ into strongly connected components $SCCs$: $\{SCC_{1},SCC_{2},\dots,SCC_{n}\}$
4:for each $SCC_{i}\in\{SCC_{1},\dots,SCC_{n}\}$ do
5:     for each vertex $v_{k}\in SCC_{i}$ do
6:         Compute in-degree score $e_{k}^{\text{in}}$
7:     end for
8:     Remove edges within $SCC_{i}$
9:     for each pair of vertices $(v_{i},v_{j})\in SCC_{i}$ do
10:         if  $e_{i}^{\text{in}}>e_{j}^{\text{in}}$  then
11:              Add edge $(v_{j}\rightarrow v_{i})$
12:         else if  $e_{i}^{\text{in}}=e_{j}^{\text{in}}$  then
13:              Add bidirectional edge $(v_{i}\leftrightarrow v_{j})$
14:         end if
15:     end for
16:end for
17:Combine all modified $SCC_{i}$ components to form a global DAG $G^{\prime}$
18:for each edge in $G^{\prime}$ do
19:     if edge is $(v_{i}\rightarrow v_{j})$ then
[…]
21:     else if edge is $(v_{i}\leftrightarrow v_{j})$ then
[…]
23:     end if
24:end for
25:Initialize empty dataset Cleaned
26:for each data point in original dataset do
27:     if preference relation matches $\mathcal{J}$ then
28:         Add data point to Cleaned
29:     end if
30:end for
31:return Cleaned
Model Helpful_Base Vicuna Oasst Koala Self-instruct
$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
Qwen2.5-Max 76.97% 0.8654 78.04% 0.9231 79.71% 0.8659 76.31% 0.8923 71.77% 0.8595
Qwen-Base 85.94% 0.9371 90.71% 0.9602 84.88% 0.9459 81.96% 0.9626 76.64% 0.8853
Qwen-Raw 79.07% 0.9244 83.21% 0.9396 77.58% 0.9066 78.66% 0.9078 73.87% 0.8853
Qwen-Cleaned (ours) 74.42% 0.8596 85.89% 0.8922 73.94% 0.8564 72.34% 0.8403 69.05% 0.8288
Table 5: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs( using the CoT Comparison (Tie Allowed) prompt template). Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data.
Model Helpful_Base Vicuna Oasst Koala Self-instruct
$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
Auto-J-13B 51.50% 0.7049 63.57% 0.7765 54.86% 0.7101 53.94% 0.7213 58.90% 0.7371
LLaMA-Base 76.41% 0.8518 66.96% 0.8461 67.40% 0.7933 69.87% 0.8180 71.03% 0.8091
LLaMA-Raw 59.03% 0.7654 60.18% 0.7718 58.21% 0.7595 57.22% 0.7457 59.95% 0.7830
LLaMA-Cleaned (ours) 40.20% 0.6523 45.36% 0.6910 43.01% 0.6288 44.38% 0.6591 41.95% 0.6423
Table 6: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model, while LLaMA-Raw, LLaMA-Random, and LLaMA-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively. For example, LLaMA-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data. Auto-J 13B is fine-tuned based on the LLaMA-2-13B-Chat model.
Model Helpful_Base Vicuna Oasst Koala Self-instruct Average
$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
Stronger LLMs
Qwen2.5-Max 63.68% 0.8047 75.36% 0.8448 64.29% 0.7883 71.52% 0.8296 64.97% 0.7803 67.96% 0.8095
Base Model
Qwen-Base 82.83% 0.9221 78.93% 0.8912 83.36% 0.9138 80.96% 0.9097 81.46% 0.9123 81.51% 0.9098
Cleaned group
Qwen-Helpful_Base-Cleaned 44.85% 0.6997 48.93% 0.7371 52.13% 0.7137 53.75% 0.7223 51.98% 0.7138 50.33% 0.7173
Qwen-Vicuna-Cleaned 47.18% 0.7395 43.93% 0.7262 52.43% 0.7276 53.48% 0.7418 57.77% 0.7483 50.96% 0.7367
Qwen-Oasst-Cleaned 45.74% 0.7040 45.00% 0.7109 46.96% 0.6942 54.85% 0.7186 51.08% 0.7227 48.73% 0.7101
Qwen-Koala-Cleaned 50.28% 0.7357 46.79% 0.7045 48.94% 0.7183 48.53% 0.7149 48.70% 0.7150 48.65% 0.7177
Qwen-Self-instruct-Cleaned 56.26% 0.7504 51.79% 0.7489 54.94% 0.7370 57.23% 0.7590 48.98% 0.6802 53.84% 0.7351
Raw group
Qwen-Helpful_Base-Raw 62.02% 0.7963 74.29% 0.8626 63.07% 0.8109 69.87% 0.8385 60.49% 0.7936 65.95% 0.8204
Qwen-Vicuna-Raw 61.13% 0.8147 57.50% 0.8034 64.59% 0.8019 60.35% 0.8150 64.23% 0.7966 61.56% 0.8063
Qwen-Oasst-Raw 60.58% 0.8099 70.89% 0.8101 55.78% 0.7734 63.83% 0.8150 61.96% 0.7746 62.61% 0.7966
Qwen-Koala-Raw 64.23% 0.8228 67.86% 0.8090 64.59% 0.7966 64.29% 0.8159 59.30% 0.7779 64.05% 0.8044
Qwen-Self-instruct-Raw 68.99% 0.8569 71.79% 0.8595 68.84% 0.8167 66.85% 0.8433 59.69% 0.7665 67.23% 0.8286
Table 7: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, Qwen-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while Qwen-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
Model Helpful_Base Vicuna Oasst Koala Self-instruct Average
$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
Stronger LLMs
Qwen2.5-Max 63.68% 0.8047 75.36% 0.8448 64.29% 0.7883 71.52% 0.8296 64.97% 0.7803 67.96% 0.8095
Base Model
Auto-J-13B 51.50% 0.7049 63.57% 0.7765 54.86% 0.7101 53.94% 0.7213 58.90% 0.7371 56.55% 0.7300
LLaMA-Base 76.41% 0.8518 66.96% 0.8461 67.40% 0.7933 69.87% 0.8180 71.03% 0.8091 70.33% 0.8237
Cleaned group
LLaMA-Helpful_Base-Cleaned 40.20% 0.6523 47.86% 0.6810 49.92% 0.6827 50.92% 0.7003 55.33% 0.7300 48.85% 0.6893
LLaMA-Vicuna-Cleaned 49.83% 0.7139 45.36% 0.6910 48.02% 0.6863 53.02% 0.7295 57.14% 0.7408 50.67% 0.7123
LLaMA-Oasst-Cleaned 50.28% 0.6655 47.68% 0.6812 43.01% 0.6288 50.73% 0.7034 54.25% 0.7056 49.19% 0.6769
LLaMA-Koala-Cleaned 44.85% 0.6588 46.61% 0.6826 44.38% 0.6591 45.51% 0.6827 50.11% 0.7031 46.29% 0.6773
LLaMA-Self-instruct-Cleaned 40.42% 0.6373 49.29% 0.6951 41.95% 0.6423 48.99% 0.6959 46.15% 0.6652 45.36% 0.6672
Raw group
LLaMA-Helpful_Base-Raw 59.03% 0.7654 60.71% 0.7888 59.42% 0.7682 61.36% 0.7965 63.78% 0.7886 60.86% 0.7815
LLaMA-Vicuna-Raw 56.92% 0.8028 60.18% 0.7718 58.81% 0.7646 60.35% 0.7943 59.69% 0.7644 59.19% 0.7797
LLaMA-Oasst-Raw 59.80% 0.7821 64.64% 0.7878 58.21% 0.7596 67.49% 0.8062 63.89% 0.7745 62.81 0.7820
LLaMA-Koala-Raw 58.03% 0.7532 53.75% 0.7601 57.22% 0.7457 57.23% 0.7597 65.59% 0.7887 58.36% 0.7615
LLaMA-Self-instruct-Raw 62.79% 0.7860 64.11% 0.7885 59.95% 0.7830 65.02% 0.8054 56.52% 0.7559 61.68% 0.7838
Table 8: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, LLaMA-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while LLaMA-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
System
You are a highly efficient assistant, who evaluates and selects the best large language model (LLMs) based on the quality of their responses to a given instruction. This process will be used to create a leaderboard reflecting the most accurate and human-preferred answers.
User
I require a leaderboard for various large language models. I’ll provide you with prompts given to these models and their corresponding outputs. Your task is to assess these responses and select the model that produces the best output from a human perspective.
## Instruction
"instruction": """{instruction}""",
## Model Outputs
Here are the unordered outputs from the models. Each output is associated with a specific model, identified by a unique model identifier.
"model_identifier": "m",
"output": """{output_1}"""
"model_identifier": "M",
"output": """{output_2}"""
## Task
Evaluate the models based on the quality and relevance of their outputs, and select the model that generated the best output. Answer by first providing a concise explanation and then end your answer by providing the model identifier of the best output. We will use the last character of your output ‘output[-1]‘ as the name of the best model, so make sure you finish with the token of the model identifiers and nothing else: ‘m‘ or ‘M‘ (no quotes, no dots, no backticks, no new lines, …). For example:
### Concise explanation …some text…
### Which is best, m or M?
M
Now is your turn.
## Your answer: "Concise explanation" followed by "Which is best, m or M?"
Table 9: The Chain-of-Thought Comparison prompt for pairwise comparison.
System
You are a highly efficient assistant, who evaluates and selects the best large language model (LLMs) based on the quality of their responses to a given instruction. This process will be used to create a leaderboard reflecting the most accurate and human-preferred answers.
User
I require a leaderboard for various large language models. I’ll provide you with prompts given to these models and their corresponding outputs. Your task is to assess these responses, and select the model that produces the best output from a human perspective. If you determine that both outputs are of equal quality or are unable to decide which one is better, you should indicate a tie by providing the identifier ‘D‘.
## Instruction
"instruction": """{instruction}""",
## Model Outputs
Here are the unordered outputs from the models. Each output is associated with a specific model, identified by a unique model identifier.
"model_identifier": "m",
"output": """{output_1}"""
"model_identifier": "M",
"output": """{output_2}"""
## Task
Evaluate the models based on the quality and relevance of their outputs, and select the model that generated the best output. Answer by first providing a concise explanation and then end your answer by providing the model identifier of the best output. If you determine that both outputs are of equal quality or cannot decide which one is better, indicate a tie by using the identifier ‘D‘. We will use the last character of your output ‘output[-1]‘ as the name of the best model, so make sure you finish with the token of the model identifiers and nothing else: ‘m‘, ‘M‘ or ‘D‘ (no quotes, no dots, no backticks, no new lines, …). For example:
### Concise explanation …some text…
### Which is best, m, M or D?
M
Now is your turn.
## Your answer: "Concise explanation" followed by "Which is best, m, M or D?"
Table 10: The Chain-of-Thought Comparison prompt (Tie Allowed) for pairwise comparison.

### v3

@@
 Abstract
-Large language models (LLMs) are widely used as evaluators for open-ended tasks, while previous research has emphasized biases in LLM evaluations, the issue of non-transitivity in pairwise comparisons remains unresolved: non-transitive preferences for pairwise comparisons, where evaluators prefer A over B, B over C, but C over A.
-Our results suggest that low-quality training data may reduce the transitivity of preferences generated by the Evaluator LLM.
-To address this, We propose a graph-theoretic framework to analyze and mitigate this problem by modeling pairwise preferences as tournament graphs.
-We quantify non-transitivity and introduce directed graph structural entropy to measure the overall clarity of preferences.
-Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max), reflecting low overall clarity of preferences.
-To address this issue, we designed a filtering strategy, ELSPR, to eliminate preference data that induces non-transitivity, retaining only consistent and transitive preference data for model fine-tuning.
-Experiments demonstrate that models fine-tuned with filtered data reduce non-transitivity by 13.78% (from 64.28% to 50.50%), decrease structural entropy by 0.0879 (from 0.8113 to 0.7234), and align more closely with human evaluators (human agreement rate improves by 0.6% and Spearman correlation increases by 0.01).
-ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction
-[…]
-[…]
+Pairwise evaluation of large language models (LLMs) has become the dominant paradigm for benchmarking open-ended tasks, yet non-transitive preferences—where evaluators prefer A over B, B over C, but C over A—fundamentally undermine ranking reliability.
+We show that this critical issue stems largely from low-quality data that contains inherently ambiguous preference pairs.
+To address this challenge, we propose ELSPR, a principled graph-theoretic framework that models pairwise preferences as tournament graphs and systematically identifies problematic training data.
+ELSPR quantifies non-transitivity through strongly connected components (SCCs) analysis and measures overall preference clarity using a novel normalized directed graph structural entropy metric.
+Our filtering methodology selectively removes preference data that induce non-transitivity while preserving transitive preferences.
+Extensive experiments on the AlpacaEval benchmark demonstrate that models fine-tuned on ELSPR-filtered data achieve substantial improvements: a 13.8% reduction in non-transitivity, a 0.088 decrease in structural entropy, and significantly enhanced discriminative power in real-world evaluation systems.
+Human validation confirms that discarded data exhibit dramatically lower inter-annotator agreement (34.4% vs.
+52.6%) and model-human consistency (51.2% vs.
+80.6%) compared to cleaned data.
+These findings establish ELSPR as an effective data self-purification approach for developing more robust, consistent, and human-aligned LLM evaluation systems.
+Code & Datasetshttps://github.com/yy0525/ELSPR
 1 Introduction
-Figure 1: Non-Transitive Preferences in LLM-as-a-Judge for Pairwise Comparisons (e.g., $A\succ B$, $B\succ C$, $C\succ A$).
-With the rapid advancement of large language model (LLM; OpenAI et al.
-2024; Qwen Team 2025; Meta AI 2024a) technology, an increasing number of models have become available, making it essential to evaluate their capabilities for selecting the most suitable one.
-However, existing benchmarks such as MMLU Hendrycks et al.
-(2020) and HELM Liang et al.
-(2022) have been shown to be insufficient for capturing performance differences in open-ended tasks Zheng et al.
-(2024).
+With the rapid advancement of large language model (LLM; openai2024gpt4technicalreport; qwen25max; llama31) technology, an increasing number of models have become available, making it essential to evaluate their capabilities for selecting the most suitable one.
+However, existing benchmarks such as MMLU (hendrycks2020measuring) and HELM (liang2022holistic) have been shown to be insufficient for capturing performance differences in open-ended tasks (zheng2024judging_mt).
 Given the lack of definitive answers in open-ended tasks, human expert evaluation is considered the gold standard.
 Yet, due to its high cost and limited scalability, the mainstream approach has shifted toward using LLM-as-a-Judge for efficient evaluation.
-Recent studies have demonstrated that powerful LLMs, such as GPT-4, can achieve high consistency with human judgments Zheng et al.
-(2024); Dubois et al.
-(2023).
-Among various approaches, pairwise comparison has emerged as the dominant paradigm, owing to its strong alignment with human preferences Samvelyan et al.
-(2024); Chen et al.
-(2024a); Li et al.
-(2023); Chiang et al.
-(2024); Liu et al.
-(2024); Liusie et al.
-(2024).
-Despite these promising results, LLM-as-a-Judge still suffers from various biases—such as position, verbosity, conformity, and self-reinforcement bias—which compromise evaluation reliability Xu et al.
-(2024b); Zheng et al.
-(2024); Koo et al.
-(2024); Ye et al.
-(2025).
-A particularly underexplored yet critical issue is the non-transitivity of preferences generated by Evaluator LLMs—e.g., $A\succ B$, $B\succ C$, $C\succ A$, where $\succ$ denotes "is preferred to", i.e., $A\succ B$ means $A$ is preferred over $B$.
+Recent studies have demonstrated that powerful LLMs, such as GPT-4, can achieve high consistency with human judgments (zheng2024judging_mt; NEURIPS2023_5fc47800).
+Among various approaches, pairwise comparison has emerged as the dominant paradigm, due to its strong alignment with human preferences (NEURIPS2024_8147a43d; chen-etal-2024-humans; alpaca_eval; chiang2024chatbot; liu2024aligning; liusie-etal-2024-llm).
+Despite these promising results, LLM-as-a-Judge still suffers from various biases—such as position, verbosity, conformity, selection, and self-reinforcement bias—which compromise evaluation reliability (xu-etal-2024-pride; zheng2024judging_mt; koo-etal-2024-benchmarking; ye2025justice; wei-etal-2024-unveiling; choi-etal-2025-mitigating).
+A particularly underexplored yet critical issue is the non-transitivity of preferences generated by evaluator LLMs (e.g., $A\succ B$, $B\succ C$, $C\succ A$), where $\succ$ denotes “is preferred to”; that is, $A\succ B$ means $A$ is preferred over $B$.
 An illustration is provided in Figure 1.
-The first study to identify this issue Xu et al.
-(2025) analyzed inconsistencies in evaluation results when GPT-4 acted as the evaluator and performed pairwise comparisons using different baseline models.
-However, the study did not propose a method to mitigate this non-transitivity.
-Alarmingly, this issue is also inherited by specialized evaluators like JudgeLM, PandaLM, and Auto-J Zhu et al.
-(2025); Wang et al.
-(2024b); Li et al.
-(2024a), which are trained via knowledge distillation from foundation models.
+Recent research (xu2025investigating) analyzed inconsistencies in the GPT-4 evaluation outcomes when conducting pairwise comparisons between different baseline models.
+Building on the widely adopted AlpacaEval framework, their findings indicate that preference non-transitivity significantly undermines the robustness of evaluation systems, leading to inconsistent model rankings under various baselines and thereby compromising the reliability of LLM-based evaluators.
+However, their analysis is limited to observing preference non-transitivity within triplets of samples and does not extend to larger sample sets.
+Moreover, they do not propose an effective method to substantially reduce the degree of non-transitivity exhibited by evaluator LLMs.
+These limitations highlight the need for more comprehensive investigations, particularly with larger sample sets, to further explore preference non-transitivity within LLM-based evaluation frameworks.
+Worryingly, this issue is also inherited by specialized evaluators such as JudgeLM, PandaLM, and Auto-J (zhu2025judgelm; wang2024pandalm; li2024generative), which are trained through the distillation of knowledge from advanced models.
 The distillation process may inadvertently propagate non-transitive judgment patterns to downstream evaluators.
 We hypothesize that the presence of low-quality training data may impair the transitivity of the preferences generated by the evaluator LLM.
-Many pairwise comparisons—particularly those from open-ended tasks—lack clear ground truth due to their subjective nature.
-Even human annotators often disagree, with inter-annotator agreement as low as 65.7% Li et al.
-(2023), indicating that such tasks inherently lack a universally accepted preference order.
-Consequently, training on ambiguous or low-quality comparisons can introduce or exacerbate non-transitive behavior in evaluators.
-This highlights the necessity of filtering unreliable data prior to model training.
-In this work, we propose ELSPR, a simple yet effective method based on graph theory to assess and mitigate preference non-transitivity in Evaluator LLMs.
-We model multi-response pairwise comparisons as tournament graphs and filter out preference data that induce the non-transitivity of overall preferences.
-To further quantify the overall clarity of preference results, we introduce the use of two-dimensional structural entropy on directed graphs.
-Our analysis reveals significant non-transitivity in advanced Evaluator LLMs (with Qwen2.5-Max exhibiting 67.96%), as well as high entropy values (0.8095 for Qwen2.5-Max).
-Experimental results show that fine-tuning models on the filtered data leads to a 13.78% reduction in preference non-transitivity, a 0.0879 decrease in structural entropy, and improved alignment with human judgments—reflected in a 0.6% increase in human agreement rate and a 0.01 gain in Spearman correlation.
+Many pairwise comparisons, particularly those from open-ended tasks, lack definitive ground truth due to their inherent subjectivity.
+Human annotators frequently demonstrate significant disagreement, with empirical studies reporting inter-annotator agreement rates as low as 65.7% (alpaca_eval), indicating that such tasks fundamentally lack universally accepted preference orderings.
+Consequently, training models with ambiguous and low-quality training data may introduce or exacerbate non-transitive preference relationships, undermining the development of stable and reliable evaluator LLMs.
+This phenomenon underscores the critical necessity of implementing robust filtering mechanisms to eliminate unreliable data points prior to model training.
+In this paper, we present a novel graph-theoretic approach to assess and mitigate preference non-transitivity in evaluator LLMs.
+Our method, ELSPR, formulates multi-response pairwise comparisons as tournament graphs and systematically filters preference data that induce to overall preference non-transitivity.
+To quantify preference clarity, we introduce a new metric based on two-dimensional structural entropy of directed graphs.
+Experimental results demonstrate that models fine-tuned on cleaned data reduce preference non-transitivity by 13.78% and structural entropy by 0.0879, while achieving more robust rankings in real-world evaluation systems, as evidenced by increased standard deviations in MT-bench evaluation framework metrics.
+Human evaluation studies further validate our approach, confirming that data inducing to overall preference non-transitivity correlates with low inter-annotator agreement among human evaluators and poor alignment between LLM evaluations and human majority votes.
+These findings underscore ELSPR’s effectiveness in developing more reliable evaluation systems for LLMs.
 Specifically, our contributions are:
-  • We introduce a graph-theoretic approach to systematically investigate the non-transitivity issue in preference data generated by Evaluator LLMs, revealing significant non-transitivity in the preference data produced by advanced LLMs.
-  • We innovatively use the two-dimensional structural entropy of the directed graph to further evaluate the overall clarity of the preferences generated by large-scale LLMs.
-  • We propose a effective filtering methodology that leverages tournament graphs to identify and remove preference data that induces non-transitivity, ensuring the retention of only consistent and transitive preferences.
-This enhances the reliability of Evaluator LLMs.
-In addition, we released our code and datasets to facilitate further research into this issue [^1] https://github.com/yy0525/ELSPR.
+  • We introduce a graph-theoretic approach to systematically study the non-transitivity problem in preference data generated by evaluator LLM, reveal significant non-transitivity, and propose a two-dimensional structural entropy to quantify preference clarity.
+  • We propose a robust filtering methodology leveraging tournament graph theory to systematically identify and eliminate preference data that induce non-transitivity in evaluator LLMs.
+Our experimental results demonstrate that this approach significantly reduces preference non-transitivity, decreases structural entropy, and substantially enhances the evaluation robustness of the resulting models when deployed in real-world evaluation systems.
+  • We empirically demonstrate through human evaluation that data causing preference non-transitivity is inherently ambiguous and low-quality, with significantly lower inter-annotator agreement (34.4% vs.
+52.6%) and lower consistency between evaluator LLM assessments and human majority votes (51.2% vs.
+80.6%).
+Figure 1: Non-Transitive Preferences in LLM-as-a-Judge for Pairwise Comparisons (e.g., $A\succ B$, $B\succ C$, $C\succ A$).
 2 Related Work
 2.1 LLM-as-a-Judge and Its Non-Transitive Preferences
-The prevailing LLM-as-a-Judge system predominantly operates through pairwise comparisons, as exemplified by frameworks such as VicunaEval, AlpacaEval, and Arena-HardChiang et al.
-(2023); Li et al.
-(2023, 2024b).
-Specifically, these systems collect responses generated by different LLMs to a given set of questions, and use advanced LLMs as judges to evaluate the preference order between two responses, assessing the relative strengths and weaknesses of different LLMs.
-Recent studies highlight that even advanced models like GPT-4 exhibit significant preference non-transitivity in LLM-as-a-Judge systems, which severely undermines the reliability of preference outcomes.
-Despite prior efforts exploring six prompt templates to alleviate preference non-transitivity, improvements have been limited Xu et al.
-(2025).
-To address this critical issue, we investigate the underlying conditions that give rise to non-transitivity in LLM-as-a-Judge systems and propose a novel method to mitigate such inconsistencies.
+The prevailing LLM-as-a-Judge paradigm operates primarily through pairwise comparisons, as evidenced in established frameworks such as VicunaEval, AlpacaEval, and Arena-Hard (vicuna2023; alpaca_eval; DBLP:journals/corr/abs-2406-11939).
+These systems collect responses generated by various LLMs for a given set of questions and then employ advanced LLMs as judges to determine preference orders between response pairs, thus evaluating the relative performance of different models.
+Recent studies demonstrate that even advanced models such as GPT-4 exhibit significant preference non-transitivity when used in evaluation systems, substantially undermining the reliability of evaluation outcomes.
+Despite systematic exploration of six distinct prompt templates, improvements in mitigating preference non-transitivity remain limited (xu2025investigating).
+This issue extends to Reinforcement Learning from Human Feedback (RLHF), where non-transitive preferences significantly impair performance.
+Recent research suggests that this non-transitivity can hinder the learning algorithm from stably converging to the global optimum, leading to preference cycles and thus hindering the effective utilization of preference data (zhou2025extragradient).
+To address this critical challenge, we systematically investigate the underlying conditions that generate non-transitivity in LLM-as-a-Judge systems and introduce a novel methodology to mitigate such inconsistencies.
 Figure 2: ELSPR (Evaluator LLM training data Self-purification non-transitive Preferences via tournament graph Reconstruction) framework overview.
-The collected preference data are modeled as a directed graph.
-Strongly Connected Components (SCCs) are identified to analyze non-transitive structures, and each SCC is reconstructed into a DAG to filter out the specific preference data responsible for the overall non-transitivity.
+(a) Raw preference data is collected via pairwise comparisons conducted by an advanced LLM.
+(b) The core analysis and filtering process: The raw data is first modeled as a tournament graph to identify cycles within SCCs.
+These cycles are then broken by reconstructing each SCC into a DAG based on in-degree ranking.
+The final global DAG serves as a rule to filter the initial raw data, separating it into a cleaned, transitively consistent training set and a discarded set of non-transitive preferences.
 2.2 Data Selection for LLM Fine-tuning
-Prior works Chen et al.
-(2024b); Ge et al.
-(2024); Li et al.
-(2024c) have highlighted the importance of filtering high-quality subsets from original datasets.
-Fine-tuning with such subsets not only enhances model performance but also reduces computational costs.
-However, this strategy remains underexplored in the context of the LLM-as-a-Judge task.
-Existing studies have primarily focused on scaling model size or improving prompt engineering, overlooking the potential benefits of training data quality.
-Our goal is to incorporate data filtering techniques into the LLM-as-a-Judge setting.
-We find that fine-tuning the LLM on preference data that has been filtered to induce non-transitivity can significantly reduce preference non-transitivity and improve overall preference clarity—even outperforming the teacher model.
+Previous research (chen2024alpagasus; ge-etal-2024-clustering; li-etal-2024-one) has established that extracting high-quality subsets from synthetic training sets is crucial for effective fine-tuning.
+This approach not only enhances the performance of the model, but also substantially reduces the computational requirements.
+However, this strategy remains underexplored in the domain of LLM-as-a-Judge.
+Existing methodologies predominantly focus on scaling model architectures or refining engineering techniques, while neglecting the potential benefits of optimizing training data quality.
 3 Methodology
-In this section, we present a graph-theoretic formulation of the LLM-as-a-Judge task using tournament graphs to analyze pairwise preferences from Evaluator LLMs.
-We then introduce quality analysis criteria for training data, including non-transitivity detection via Strongly Connected Components (SCCs) and preference clarity analysis via graph entropy.
-Finally, we propose a data filtering method to mitigate non-transitive preferences.
+This section presents a graph-theoretic analysis method using tournament graphs to analyze pairwise preferences from evaluator LLMs.
+The method introduces quality analysis criteria for training data, including non-transitivity detection via strongly connected components (SCCs) and preference clarity analysis via graph entropy.
+Additionally, a data filtering method is proposed to mitigate non-transitive preferences.
 Figure 2 outlines the overall framework, with implementation details provided in the following subsections.
 3.1 Background
-The typical LLM evaluation systems rely on the transitivity assumption, that is, if the evaluation results are $A\succ B$ and $B\succ C$, then it is assumed that $A\succ C$ However, evaluations in the real world often reveal non-transitivity preference cycles ($A\succ B\succ C\succ A$), which renders traditional ranking methods ineffective Xu et al.
-(2025).
-In order to overcome this fundamental limitation, we propose a tournament-based evaluation framework grounded in directed graphs, which facilitates the application of graph theory methods for a deeper analysis of the degree of non-transitivity in preferences and the overall clarity of the preferences.
 Modeling preferences generated by LLM as a tournament graph.
-For each question $q_{i}\in Q$, given the response set $A_{i}=\{a_{1},a_{2},...,a_{n}\}$ from $n$ LLMs, we construct a tournament graph $G_{i}=(V_{i},E_{i})$ through the following procedure:
-vertices: set $V_{i}=\{v_{1},v_{2},...,v_{n}\}$ corresponds to responses $A_{i}=\{a_{1},a_{2},...,a_{n}\}$.
-Edges: Defined by preferences generated by LLM:
-$E_{i}=\bigcup_{\begin{subarray}{c}1\leq j,k\leq n\\ j\neq k\end{subarray}}\begin{cases}v_{k}\rightarrow v_{j},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`win'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`lose'}\end{aligned}\\[10.00002pt] v_{j}\rightarrow v_{k},&\begin{aligned} \text{if }\mathcal{J}(a_{j},a_{k})=\text{`lose'}\\ \text{ and }\mathcal{J}(a_{k},a_{j})=\text{`win'}\end{aligned}\\[10.00002pt] v_{j}\leftrightarrow v_{k},&\text{otherwise}\end{cases}$ (1)
+For each question $q_{i}\in Q$, given the response set $A_{i}=\{a_{1},a_{2},...,a_{n}\}$ from $n$ LLMs, we construct a tournament graph $G_{i}=(V_{i},E_{i})$ through the following procedure: vertices: set $V_{i}=\{v_{1},v_{2},...,v_{n}\}$ corresponds to responses $A_{i}=\{a_{1},a_{2},...,a_{n}\}$.
+Edges: Defined by preferences generated by evaluator LLM:
 […]
-Considering the common position bias in Evaluator LLMs Wang et al.
-(2024a), we apply position swapping by comparing each response pair in both orders: $\mathcal{J}(a_{j},a_{k})$ and $\mathcal{J}(a_{k},a_{j})$.
-This ensures a more robust and balanced assessmentZheng et al.
-(2024).
+Here, $\mathcal{J}(a_{j},a_{k})\in\{\text{`win'},\text{`lose'}\}$ denotes the pairwise comparison result between answers $a_{j}$ and $a_{k}$, where $a_{j}$ appears before $a_{k}$ in the prompt.
+Considering the common position bias in evaluator LLMs (wang-etal-2024-large-language-models-fair), we apply position swapping by comparing each response pair in both orders: $\mathcal{J}(a_{j},a_{k})$ and $\mathcal{J}(a_{k},a_{j})$.
+This ensures a more robust and balanced assessment (zheng2024judging_mt).
 If the preferences differ across orders—indicating possible position bias—a bidirectional edge is added between the corresponding vertices to represent a ‘tie’.
-3.2 Quality Analysis Criteria for LLM-as-a-Judge Training Data
-Based on the directed graph representation of the tournament we constructed above (Section 3.1).
-We propose a quality analysis criteria for the training data of Evaluator LLM, which utilizes SCCs analysis to measure the degree of preference non-transitivity and employs two-dimensional structural entropy of directed graph structure entropy to evaluate the overall clarity of preferences.
-3.2.1 Analysis of Non-Transitivity in Preferences Based on SCCs
-The SCCs of a directed graph are maximal subgraphs in which, for any pair of vertices, there exists a directed path from one to the other and vice versa.
-This mutual reachability property aligns with the notion of preference cycles, which directly reflects non-transitivity in the preferences of the Evaluator LLM.
-A directed path from vertex $v_{i}$ to $v_{j}$ can be interpreted as a preference chain $a_{j}\succ\cdot\succ a_{i}$, while a reverse path from $v_{j}$ to $v_{i}$ corresponds to $a_{i}\succ\cdot\succ a_{j}$.
-The presence of both paths implies conflicting preferences and thus a violation of transitivity.
-We use Tarjan algorithm to identify the SCCs of a directed graph.
-Tarjan algorithmTarjan (1972) is an efficient method for finding SCCs, based on depth-first search (DFS), with a time complexity of $O(N+M)$, where $N$ is the number of vertices and $M$ is the number of edges.
-Furthermore, since non-transitive relations occur between three or more vertices, we focus on components with a size greater than 2.
-Additionally, we filter out SCCs where there exists a bidirectional edge between every pair of vertices to eliminate cases where the relationship between vertices is a ‘tie’.
-For example, in the case of $(A=B),(B=C),(C=A)$, although $(A,B,C)$ forms an SCC, the preference relation is transitive.
-We denote the set of all SCCs in the directed graph $G$ as $SCCs(G)$,the following formula to obtain SCCs containing vertices with non-transitive relationships:
-$\mathcal{S}_{\text{n-t}}=\left\{S\in SCCs(G)\,\middle|\,\begin{aligned} |S|>2\land\exists v_{j},v_{k}\in S,\\ (v_{j}\leftrightarrow v_{k})\notin E(S)\end{aligned}\right\}.$ (2)
-To quantify the degree of non-transitivity in the preferences generated by the Evaluator LLM, we compute the ratio of non-transitive vertices to the total number of vertices across all directed graphs:
+3.2 Quality Analysis Framework for Evaluator LLM Training Data
+We introduce a framework for analyzing evaluator LLM training data quality.
+Leveraging directed graph representations of preference tournaments, the framework introduces two metrics: SCC analysis quantifying preference non-transitivity and two-dimensional structural entropy measuring preference clarity.
+Quantifying preference non-transitivity via SCC analysis.
+SCCs are maximal subgraphs where any two vertices are mutually reachable through directed paths.
+This property provides a natural mechanism for identifying preference cycles—a direct manifestation of non-transitivity in evaluator LLM judgments.
+When vertex $v_{i}$ can reach vertex $v_{j}$ via a directed path, this indicates a preference chain $a_{j}\succ\cdot\succ a_{i}$.
+The simultaneous existence of paths from $v_{i}$ to $v_{j}$ and from $v_{j}$ to $v_{i}$ represents contradictory preference relationships, signaling a violation of transitivity.
+We employ Tarjan’s algorithm (doi:10.1137/0201010) to efficiently identify SCCs in our directed graphs.
+Tarjan’s algorithm operates in linear time, specifically $\mathcal{O}(|V|+|E|)$, where $|V|$ and $|E|$ denote the number of vertices and edges, respectively, ensuring scalability to large graphs.
+Since non-transitivity requires at least three elements in a cycle, we focus on SCCs containing more than two vertices.
+Additionally, we exclude cases where every vertex pair shares bidirectional edges (indicating ‘tie’), as these represent consistent preference relations.
+For example, the pattern $(A=B)$, $(B=C)$, $(C=A)$ constitutes a valid SCC but maintains transitive preference relationships.
+Formally, we identify the set of non-transitive SCCs as:
+$\mathcal{S}_{\text{n-t}}=\left\{S\in\text{SCCs}(G)\,\middle|\,\begin{aligned} |S|>2\land\exists v_{j},v_{k}\in S,\\ (v_{j}\leftrightarrow v_{k})\notin E(S)\end{aligned}\right\},$ (2)
+where $\text{SCCs}(G)$ represents all SCC in graph $G$, $|S|$ denotes the component size, and $(v_{j}\leftrightarrow v_{k})\notin E(S)$ indicates the absence of bidirectional edges between some vertices.
+To quantify the prevalence of non-transitivity across our entire training set, we compute the non-transitivity ratio:
 $\rho_{\text{non-trans}}=\frac{\sum_{q_{i}\in\mathbf{Q}}|S_{\text{n-t}}(G_{i})|}{\sum_{q_{i}\in\mathbf{Q}}|V_{i}|},$ (3)
-where the numerator represents the total number of vertices belonging to non-transitive SCCs across all questions $\mathbf{Q}$, and the denominator represents the total number of vertices across all graphs.
-A higher value of $\rho_{\text{non-trans}}$ indicates a greater prevalence of non-transitive preferences.
-3.2.2 Analysis of Preference Clarity Based the Entropy of Directed Graph Structures
-The Shannon entropyShannon (1948) is a fundamental metric in information theory used to quantify the uncertainty associated with a random variable.
-It represents the expected level of uncertainty in a system: the greater the uncertainty, the higher the entropy.
-It is defined as:
-$H(X)=-\sum p(x)\log p(x),$ (4)
-where $p(x)$ denotes the probability distribution of the event $x$.
-Structural entropy Li and Pan (2016) extends Shannon entropy to directed graphs, providing a measure of system uncertainty and the complexity of relationships within the graph.
-These concepts have been widely applied across various domains Zou et al.
-(2024); Duan et al.
-(2024); Peng et al.
-(2024); Hou et al.
-(2025).
-We introduce the Two-Dimensional Structural Entropy of a Directed Graph to analyze the overall clarity of preferences in Evaluator LLM.
-Two-Dimensional Structural Entropy of a Directed Graph: The core idea of two-dimensional structural entropy is to partition the graph’s vertices into distinct communities and compute the complexity of both intra-community and inter-community relationships.
-This framework effectively quantifies the overall structural uncertainty and the complexity of relationships within the graph.
-In this work, we innovatively adopt SCCs as the fundamental communities, motivated by properties from directed graph theory.
-Within an SCC, cyclic and non-transitive preference relations exist, while between SCCs, the preference relations are strictly transitive.
-This distinction allows us to quantitatively assess the overall preference clarity by computing the two-dimensional structural entropy of the tournament graphs constructed by Evaluator LLM.
-It is important to note that SCCs composed of a single vertex inherently exhibit strict transitivity.
-The preference relations among such single-vertex components form a linear order that does not increase the complexity of the preference results.
-For example, if the original graph contains a preference $A\succ B$, the overall preference is clear.
-When new preferences $B\succ C$ and $A\succ C$ are added, the resulting linear order $A\succ B\succ C$ still remains clear and consistent — since the vertices $A$, $B$, and $C$ each form their own singleton SCC.
+where the numerator represents the total number of vertices in non-transitive SCCs across all questions $\mathbf{Q}$, and the denominator represents the total number of vertices across all graphs.
+This metric ranges from 0 to 1, with higher values indicating greater prevalence of non-transitive in the evaluator LLM’s preferences.
+A perfectly transitive preference system would yield $\rho_{\text{non-trans}}=0$, while completely cyclical preferences would approach $\rho_{\text{non-trans}}=1$.
+Analysis of preference clarity based the structural entropy of directed graph.
+While the SCC-based approach effectively quantifies preference non-transitivity, this metric alone is insufficient to characterize preference linearity.
+For instance, a dataset comprised entirely of preference ‘tie’ would exhibit perfect transitivity yet fail to establish any linear ordering.
+To address this limitation, we introduce directed graph entropy as a complementary measurement.
+Structural entropy (7456290) extends Shannon entropy (6773024) to directed graphs, providing a measure of system uncertainty and the complexity of relationships within the graph.
+These concepts have been widely applied across various domains  (10.1145/3616855.3635820; Duan_Chen_Liu_Liu_Yue_Li_2024; 10.1145/3660522; Hou_Zhu_Liu_Su_Xia_Wu_Xu_2025).
+We introduce and refine the two-dimensional structural entropy metric for directed graphs to analyze the overall clarity of preferences in evaluator LLMs.
+A crucial observation drives our methodology is that SCCs composed of single vertices inherently exhibit strict transitivity.
+Preference relations among such singleton components naturally form a linear order that does not increase preference complexity.
+For example, consider preferences $A\succ B$, $B\succ C$, and $A\succ C$.
+These form a clear linear order $A\succ B\succ C$, where each vertex constitutes its own singleton SCC.
 This clarity reflects the inherent transitivity of singleton SCCs.
-Based on this observation, we exclude interactions among pure singleton SCCs when computing structural entropy.
-Instead, we retain only two types of interactions: (1) between singleton SCCs and multi-vertice SCCs, and (2) between multi-vertice SCCs.
-This selective inclusion ensures a more accurate and meaningful quantification of global structural complexity.
-Consequently, the closer the priority relations among vertices are to a linear order, the lower the resulting entropy.
-The definition of the two-dimensional structural entropy of a directed graph based on SCC decomposition is as follows: Given a directed graph $G=(V,E)$, $n=|V|$ denote the number of vertices.
-Based on the SCC decomposition $SCCs(G)=\{SCC_{1},\dots,SCC_{L}\}$.
-For a vertex $v\in V$, we use $d_{\text{in}}(v)$ and $d_{\text{out}}(v)$ to represent the in-degree and out-degree of $v$ in $G$, respectively.
-For any $SCC_{i}\subseteq SCCs$, define its volume as the total in-degree of its vertices:
-$v(SCC_{i})=\sum_{v\in SCC_{i}}d_{\text{in}}(v).$ (5)
-Accordingly, $v(G)$ denotes the total in-degree of the entire directed graph.
-The two-dimensional structural entropy computation formula is as follows:
-$H^{2}(G)=-\sum_{j=1}^{L}\frac{g_{j}}{v(G)}\log_{2}\frac{v(SCC_{j})}{v(G)}-\sum_{j=1}^{L}$ (6)
-$\frac{v(SCC_{j})}{v(G)}\left(\sum_{v\in SCC_{j}}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\log_{2}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\right),$
-where, $g_{j}$ denotes the total number of incoming edges to $V_{j}$ from external vertices.
-Specifically, it counts the edges between singleton and multi-vertice SCCs, as well as those among multi-vertice SCCs.
-It serves as a measure of the external influence on $V_{j}$.
-Specifically, the normalized structural entropy for a graph $G$ is defined as:
-$\tau(G)=\frac{H^{2}(G)}{\log n},$ (7)
-where $H^{2}(G)$ represents the two-dimensional structural entropy of $G$, and $n$ denotes the number of vertices in the graph.
-To evaluate the overall preference clarity of the Evaluator LLM across the set of questions $Q$, we compute the average normalized structural entropy, defined as follows:
+Based on this insight, we adopt SCCs as the basic community units when calculating structural entropy.
+To ensure accurate quantification of meaningful structural complexity, we exclude interactions between pure single-point SCCs.
+Instead, we retain only interactions between singleton SCCs and multi-vertex SCCs, as well as interactions between different multi-vertex SCCs.
+This selective approach ensures more accurate quantification of meaningful structural complexity.
+Consequently, preference relations closer to a linear order produce lower entropy values.
+Figure 3: Cases of High and Low Structural Entropy in Preference Tournaments.
+Figure 3 illustrates two contrasting scenarios that provide an intuitive explanation of the two-dimensional structural entropy in tournament graph preference relations.
+Figure 3 (a) depicts a high entropy case with a chaotic tournament containing multiple preference cycles resulting in complex non-transitive relationships.
+In contrast, Figure 3 (b) shows a low entropy case with a perfectly ordered tournament where responses form a clear linear hierarchy, producing a simple transitive structure.
+Given a directed graph $G=(V,E)$ with $n=|V|$ vertices, we perform SCC decomposition resulting in $\text{SCCs}(G)=\{SCC_{1},SCC_{2},\ldots,SCC_{L}\}$, where $L$ is the total number of SCCs.
+For any vertex $v\in V$, we denote $d_{\text{in}}(v)$ and $d_{\text{out}}(v)$ as the in-degree and out-degree of $v$ in $G$.
+For any $SCC_{i}$, we define its volume as $v(SCC_{i})=\sum_{v\in SCC_{i}}d_{\text{in}}(v)$, while $v(G)$ represents the total in-degree of the entire graph.
+The two-dimensional structural entropy is computed as:
+$H^{2}(G)$ $=-\sum_{j=1}^{L}\frac{g_{j}}{v(G)}\log_{2}\frac{v(SCC_{j})}{v(G)}$
+$-\sum_{j=1}^{L}\frac{v(SCC_{j})}{v(G)}\left(\sum_{v\in SCC_{j}}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\log_{2}\frac{d_{\text{in}}(v)}{v(SCC_{j})}\right),$ (4)
+The first term captures the entropy of the partition (inter-community complexity), while the second term captures the weighted average of entropies within each SCC (intra-community complexity).
+The variable $g_{j}$ represents the number of incoming edges to $SCC_{j}$ from vertices outside this component, specifically counting edges between singleton SCCs and multi-vertex SCCs, as well as edges between different multi-vertex SCCs.
+This value quantifies the external influence on each SCC and plays a crucial role in measuring cross-community complexity.
+To facilitate meaningful comparisons across graphs of different sizes, we normalize the structural entropy. The normalized structural entropy for a graph $G$ is defined as:
+$\tau(G)=\frac{H^{2}(G)}{\log_{2}n}.$ (5)
+Normalization is necessary since raw entropy increases with graph size.
+Dividing by $\log_{2}n$ scales the entropy to a range between 0 and 1, where values closer to 0 indicate more ordered structures approaching linear hierarchy, while values closer to 1 suggest higher complexity and less clear preference ordering.
+The upper bound $\log_{2}n$ represents the maximum possible entropy for a graph with $n$ vertices, corresponding to a completely uniform distribution of structural information.
+To evaluate the overall preference clarity of an evaluator LLM across a set of questions $\mathbf{Q}$, we compute the average normalized structural entropy:
 $\tau_{\text{avg}}=\frac{\sum_{q_{i}\in\mathbf{Q}}\tau(G_{i})}{|\mathbf{Q}|},$ (6)
-where $\tau(G_{i})$ is the normalized structural entropy for the graph $G_{i}$ corresponding to question $q_{i}$, and $|\mathbf{Q}|$ is the total number of questions in the set $\mathbf{Q}$.
-A higher value of $\tau_{\text{avg}}$ signifies that the preference relations generated by the Evaluator LLM are more complex, indicating a lower level of overall preference clarity.
-3.3 Filtering Strategy for Preference Data That Induces Non-Transitivity
-After decomposing a cyclic directed graph into multiple SCCs, if the cycles within each SCC can be eliminated and each SCC can be reconstructed as a DAG, then the entire directed graph can likewise be transformed into a Directed Acyclic Graph (DAG).
-Building upon this observation, our framework transforms cyclic directed graphs into DAGs by converting each SCC into a DAG, preserving inter-SCC preference relations.
-This process ensures the global graph structure becomes acyclic without altering inter-component preferences.
-Subsequently, we derive transitive preference relations from the final DAG to filter Evaluator LLMs’ preference data, facilitating self-purification of the training data.
-The detailed procedure is as follows:
+Here, $\tau(G_{i})$ is the normalized structural entropy for the graph $G_{i}$ corresponding to question $q_{i}$, and $|\mathbf{Q}|$ is the total number of questions in the set $\mathbf{Q}$.
+The $\tau_{\text{avg}}$ metric quantifies preference clarity: higher values indicate complex, non-transitive patterns, while lower values reflect relationships closer to linear order.
+It enables objective assessment of consistency across evaluator LLMs, where lower entropy denotes more coherent and interpretable preferences.
+3.3 Filtering Strategy for Preference Data That Induce Non-Transitivity
+If cycles within each SCC of a directed graph can be eliminated to form DAGs, the entire graph can be transformed into a DAG.
+Our framework leverages this property by converting each SCC into a DAG while preserving inter-SCC preference relations, ensuring global acyclicity without altering component relationships.
+From the resulting DAG, we derive transitive preference relations to filter evaluator LLMs’ preference data, enabling training data self-purification.
+The procedure is:
   1. 1.
@@
   2. 2.
-[…]
+    After processing all SCCs, the original cyclic graph is transformed into a DAG.
+This DAG is used to filter the training data as follows: For bidirectional edges $(v_{i}\leftrightarrow v_{j})$, the correct preferences are recorded as $\mathcal{J}(a_{i},a_{j})=\text{`tie'}$ and $\mathcal{J}(a_{j},a_{i})=\text{`tie'}$.
+For unidirectional edges $(v_{i}\rightarrow v_{j})$, the correct preferences are recorded as $\mathcal{J}(a_{i},a_{j})=\text{`lose'}$ and $\mathcal{J}(a_{j},a_{i})=\text{`win'}$.
   3. 3.
-    The training data is traversed sequentially, and all data consistent with the correct preferences are added to the dataset "Cleaned".
-See Appendix E for detailed algorithm information.
+    The training data are traversed sequentially.
+All instances that are consistent with the correct preferences are added to the training set “Cleaned”, while the remaining data are placed in the training set “Discarded”.
+The final “Cleaned” training data only retains linearly transitive preference data, resulting in $\rho_{\text{non-trans}}=0$ and $\tau_{\text{avg}}=0$.
+Our approach balances optimality and scalability, with an overall time complexity of $\mathcal{O}(|V|^{2}+|E|)$, consisting of $\mathcal{O}(|V|^{2})$ for graph construction, $\mathcal{O}(|V|+|E|)$ for SCC decomposition, $\mathcal{O}(|V|^{2})$ for intra-SCC reconstruction, and $\mathcal{O}(|E|)$ for edge filtering.
+The detailed algorithmic procedure is provided in Appendix A.
 4 Experiment Setup
-Figure 3: Comparison of data volumes between "Raw" and "Cleaned" training sets across datasets. The "Cleaned" dataset’s volume is approximately 80% of the "Raw" dataset for each dataset.
 Model Helpful_Base Vicuna Oasst Koala Self-instruct
@@
 Stronger LLMs
-Qwen2.5-Max 63.68% 0.8047 75.36% 0.8448 64.29% 0.7883 71.52% 0.8296 64.97% 0.7803
+Qwen2.5-Max 63.7% 0.805 75.4% 0.845 64.3% 0.788 71.5% 0.830 65.0% 0.780
 Base models and Variants
-Qwen-Base 82.83% 0.9221 78.93% 0.8912 83.36% 0.9138 80.96% 0.9097 81.46% 0.9123
-Qwen-Raw 62.02% 0.7963 57.50% 0.8034 55.78% 0.7734 64.29% 0.8159 59.69% 0.7665
-Qwen-Random 60.58% 0.7899 63.57% 0.8078 58.81% 0.7711 60.35% 0.8116 57.14% 0.7644
-Qwen-Cleaned (ours) 44.85% 0.6997 43.93% 0.7262 46.96% 0.6942 48.53% 0.7149 48.98% 0.6802
+Qwen-Base 82.8% 0.922 78.9% 0.891 83.4% 0.914 81.0% 0.910 81.5% 0.912
+Qwen-Raw 62.0% 0.796 57.5% 0.803 55.8% 0.773 64.3% 0.816 59.7% 0.767
+Qwen-Random 60.6% 0.790 63.6% 0.808 58.8% 0.771 60.4% 0.812 57.1% 0.764
+Qwen-Cleaned (ours) 44.9% 0.700 43.9% 0.726 47.0% 0.694 48.5% 0.715 49.0% 0.680
+LLaMA-Base 76.4% 0.852 67.0% 0.846 67.4% 0.793 69.9% 0.818 71.0% 0.809
+LLaMA-Raw 59.0% 0.765 60.2% 0.772 58.2% 0.760 57.2% 0.746 60.0% 0.783
+LLaMA-Cleaned (ours) 40.2% 0.652 45.4% 0.691 43.0% 0.629 44.4% 0.659 42.0% 0.642
 Table 1: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs.
-Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw", "Random", and "Cleaned" training sets, respectively.
-For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base dataset.
-Dataset: In this study, we conduct experimental validation using the AlpacaEval benchmark (Li et al., 2023).
-AlpacaEval is specifically designed to assess the overall capabilities of LLMs in open-domain, instruction-following tasks, covering a wide range of evaluation scenarios such as reasoning and text generation.
-The benchmark comprises five datasets: Helpful_Base, Oasst, Koala, Vicuna, and Self-Instruct.
-Teacher Model Selection: We evaluated the performance of Qwen2.5-MaxQwen Team (2025) using the 2.5k human-annotated samples provided by AlpacaEval, employing the Chain-of-Thought Comparison (CoT Comparison) Li et al.
-(2023) .
-The detailed prompt template is provided in Appendix C.1.
-Qwen2.5-Max demonstrated exceptional performance in both human agreement rate, achieving an agreement level of 68.9%.
-Based on these results, we selected Qwen2.5-Max as our teacher model.
-Experiment Details: We selected 21 representative LLMs from the AlpacaEval leaderboard and divided them into two groups based on the experimental design: 14 LLMs were used to collect preference data for constructing the training set, while the remaining 7 LLMs were employed to build the testing set for performance validation(see Appendix A for details).
-Throughout the experiment, we consistently applied the CoT Comparison prompt template for response generation, with the decoding temperature fixed at 0.
-Preference data for constructing the training set was generated by invoking the official Qwen2.5-Max API.
-The unfiltered Raw dataset was labeled as "Raw", and subsequently filtered according to the steps described in Section 3.3, resulting in the "Cleaned" dataset.
-The data distribution across datasets is illustrated in Figure 3.
-For comparative analysis, we fine-tuned models using both the "Raw" and "Cleaned" training sets derived from five datasets, ultimately obtaining a total of 10 fine-tuned models.
-In the experiments, we used Qwen2.5-7B-Instruct as the base model and fine-tuned it using the Low-Rank Adaptation (LoRA) methodHu et al.
-(2022), with a rank of 8.
-The model was trained for 3 epochs, with a learning rate of $1\times 10^{-4}$ and a batch size of 16.
+Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the “Raw”, “Random”, and “Cleaned” training sets, respectively.
+For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the ”Cleaned” training set derived from filtered Helpful_Base training set.
+4.1 Dataset
+In this study, we conduct experimental validation using the AlpacaEval benchmark (alpaca_eval).
+AlpacaEval is specifically designed to assess the overall capabilities of LLMs in open-ended tasks, covering a wide range of evaluation scenarios such as reasoning and text and code generation.
+The benchmark comprises five datasets, Helpful_Base, Oasst, Koala, Vicuna, and Self-Instruct.
+4.2 Preference Data Collection
+We evaluated Qwen2.5-Max (qwen25max) on 2.5k human-annotated samples from AlpacaEval using the chain-of-thought (COT) comparison (alpaca_eval).
+Qwen2.5-Max achieved a human agreement rate of 68.9%.
+Based on these results, we selected it as our teacher model for preference data generation.
+4.3 Experiment Details
+21 representative LLMs were selected from the AlpacaEval leaderboard (14 for training, 7 for testing).
+All experiments employed CoT comparison templates with temperature set to 0, and preferences were generated via the Qwen2.5-Max API.
+The unfiltered “Raw” datasets were processed according to the procedure described in Section 3.3 to produce “Cleaned” training sets; the corresponding distributions are illustrated in Figure 4.
+Qwen2.5-7B-Instruct was fine-tuned on both versions of each of the five datasets using LoRA (hu2022lora) (rank =8=8, 3 epochs, learning rate $1\times 10^{-4}$, batch size =16=16).
+Figure 4: Comparison of data volumes between “Raw” and “Cleaned” training sets across datasets. The “Cleaned” training set’s volume is approximately 80% of the “Raw” training set for each dataset.
 5 Results and Analysis
-In this section, we validate the effectiveness of the proposed filtering method on five datasets.
-We utilize $\rho_{\text{non-trans}}$, as defined in Equation 3, to quantify the degree of preference non-transitivity in the evaluation of LLMs.
-A higher value of $\rho_{\text{non-trans}}$ indicates a greater extent of non-transitive preferences.
-Additionally, we employ $\tau_{\text{avg}}$, as defined in Equation 8, to assess the overall clarity of preferences, where a lower value reflects improved preference clarity.
-Furthermore, we provide a comprehensive measure of the alignment between model preferences and human preferences by calculating the human agreement rate and the spearman correlation coefficient between the model’s evaluations and human assessments.
-Additionally, we investigate two scenarios in which Evaluator LLMs are prone to exhibit non-transitive preferences, see Appendix D.
+This section verifies the effectiveness of the proposed filtering method on five datasets.
+The analysis begins with an examination of preference non-transitivity degree and overall preference clarity of evaluator LLMs.
+To further demonstrate the method’s effectiveness, experiments based on actual usage scenarios are conducted using the MT-bench evaluation framework to calculate the standard deviation between final results.
+Additionally, a detailed data quality analysis examines two key aspects of manual evaluation: (1) consistency between manual annotators, and (2) consistency between evaluation results and human majority vote.
 5.1 Main Results
-5.1.1 Analysis of Non-Transitivity and Clarity in Preferences
-As shown in Table 1, across all testing sets, models fine-tuned on the "Cleaned" training set exhibit the lowest degree of preference non-transitivity and the highest clarity of preferences, surpassing even the teacher model Qwen2.5-Max.
-This highlights the effectiveness of our proposed data filtering method.
-Additionally, models fine-tuned on the "Raw" dataset perform comparably to the teacher model Qwen2.5-Max, demonstrating that data distillation can indeed enhance a model’s evaluation capability relative to its original performance.
-However, due to the inherent preference non-transitivity of the teacher model, direct distillation may learn low-quality preference data, thereby limiting the ability of the distilled model to surpass the teacher model and resulting in similar performance.
-To validate that models fine-tuned on the "Cleaned" training set achieve lower preference non-transitivity and higher preference clarity even on "unseen" questions, we performed cross-validation by testing each fine-tuned model on five testing sets.
-As shown in Table 2, models fine-tuned on Cleaned data show a 13.78% average reduction in non-transitivity and a 0.0879 decrease in normalized structural entropy compared to those trained on Raw data, indicating clearer overall preferences.
-Detailed results are in Appendix B.3.
-Notably, even advanced Evaluator LLMs (e.g., Qwen2.5-Max at 67.96%) exhibit high non-transitivity and entropy (0.8591 and 0.8095, respectively), reflecting limited preference clarity.
-Training Set $\rho_{\text{non-trans}}$ $\tau_{\text{avg}}$
-Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
-Helpful_Base 65.95% 50.33% -15.62% 0.8204 0.7173 -0.1031
-Vicuna 61.56% 50.96% -10.60% 0.8063 0.7367 -0.0696
-Oasst 62.61% 48.73% -13.88% 0.7966 0.7101 -0.0865
-Koala 64.05% 48.65% -15.40% 0.8044 0.7177 -0.0867
-Self-instruct 67.23% 53.84% -13.39% 0.8286 0.7351 -0.0935
-Average 64.28% 50.50% -13.78% 0.8113 0.7234 -0.0879
-Table 2: Comparison of average preference non-transitivity and average clarity between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct across testing sets.
-The leftmost column indicates the source dataset used for training, while other columns show model average performance on five testing sets.
-For example, the "Raw" column under Helpful_Base shows results for unfiltered data, and the "Cleaned" column reflects filtered data.
-5.1.2 Human Agreement Rate and Spearman Correlation Analysis.
-We evaluate the alignment of fine-tuned Evaluator LLMs with 2.5k human-annotated preference labels from the AlpacaEval dataset, which contains around 650 questions with two candidate responses each, annotated by four human evaluators Li et al.
-(2023).
-The questions are sourced from Helpful_Base, Oasst, Koala, Vicuna, Self-Instruct, and Alpaca.
-Human Agreement Rate: This metric quantifies the alignment between Evaluator LLMs and human annotators’ majority preferences.
-For each instruction, four predictions are generated by the Evaluator LLM.
-Using a leave-one-out approach, each prediction is compared against the majority vote of the remaining three human annotations, excluding the corresponding human annotation.
-The final score is computed as the average accuracy across all four comparisons per instruction.
-Spearman Correlation:This metric evaluates the Spearman correlation between rankings based on Evaluate LLMs’ preferences and those derived from human preferences.
-As shown in Table 3, models fine-tuned on the corresponding "Cleaned" training sets of Vicuna, Oasst, koala and Self-instruct exhibit higher human agreement rate compared to those fine-tuned on the "Raw" training sets, with improvements ranging from 0.6% to 1.2%.
-Additionally, models fine-tuned on the corresponding "Cleaned" training sets of Helpful_Base, Vicuna, and Self-instruct exhibit higher Spearman correlation compared to those fine-tuned on the "Raw" training sets.
-Notably, on the Helpful_Base "Cleaned" training set achieves the largest improvement, with the Spearman correlation increasing from 0.93 to 0.97.
-5.2 Ablation Experiments
-Ablation of Different Base Models: We conducted repeated experiments using the LLaMA3.1-8B-Instruct model as the base model.
-As shown in Appendix B.2, models trained on the "Cleaned" training set outperform those trained on the "Raw" training set and also surpass the advanced open-source Evaluator LLM Auto-J 13B.
-We further validated that this superiority holds true even for "unseen" questions, with detailed results provided in Appendix B.3, further highlighting the importance of high-quality data in enhancing model performance.
-Ablation of Different Data Filtering Methods: As shown in Figure 3, the "Cleaned" training set constitutes approximately 80% of the "Raw" training set.
-To validate the effectiveness of our data filtering method, we randomly filtered 20% of the training data and fine-tuned the Qwen2.5-7B-Instruct model using the remaining 80%, denoted as Qwen-Random.
-As shown in Table 1, the overall performance is similar to the model trained using the "Raw" training set, and showed higher preference non-transitivity on the Vicuna testing set.
-This comparison underscores the effectiveness of our targeted filtering strategy in enhancing data quality.
-Ablation of Different Prompt Forms: We conducted additional experiments using a prompt format that includes ‘tie’ preference outcomes.
-The detailed prompt template is provided in Appendix C.2, and the experimental setup and results are presented in Appendix B.1, further validating the effectiveness of our data filtering method.
-Notably, when using prompts that include ‘tie’ preference outcomes, Evaluate LLMs exhibit higher preference non-transitivity and lower overall preference clarity compared to prompts without ‘tie’ outcomes.
-This finding is consistent with the results reported in prior work Xu et al.
-(2025).
-Training Set Human Agreement Rate Spearman Correlation
-Raw Cleaned $\Delta$ Raw Cleaned $\Delta$
-Helpful_Base 66.9% 66.9% +0.0% 0.93 0.97 +0.04
-Vicuna 65.2% 65.8% +0.6% 0.97 0.98 +0.01
-Oasst 66.4% 67.6% +1.2% 0.93 0.93 +0.00
-Koala 66.4% 66.9% +0.5% 0.98 0.98 +0.00
-Self-instruct 67.8% 68.3% +0.5% 0.98 1.00 +0.02
-Average 66.5% 67.1% +0.6% 0.96 0.97 +0.01
-Table 3: Comparison of Human Agreement Rate and Spearman Correlation between models fine-tuned on "Cleaned" and "Raw" datasets using Qwen2.5-7B-Instruct. The structure matches Table 2.
+Analysis of preference non-transitivity and clarity.
+As shown in Table 1, models fine-tuned on the “Cleaned” training set demonstrate the lowest preference non-transitivity and highest preference clarity across all datasets, including Qwen2.5-Max, highlighting the effectiveness of our proposed data filtering methodology.
+To verify that models fine-tuned on the “Cleaned” training set achieve lower preference non-transitivity and higher preference clarity even on “unseen” questions, we conducted cross-validation as our out-of-domain experiment by testing each fine-tuned model across five test sets.
+As demonstrated in Table 2, models fine-tuned on “Cleaned” training set exhibited an average reduction in non-transitivity of 13.8% compared to those trained on “Raw” training set.
+Table 3 shows a reduction of 0.088 in normalized structural entropy, indicating clearer overall preferences.
+Detailed results are in Appendix B.
+Dataset Raw Cleaned $\Delta$
+Helpful_Base 66.0% 50.3% -15.7%
+Vicuna 61.6% 51.0% -10.6%
+Oasst 62.6% 48.7% -13.9%
+Koala 64.1% 48.7% -15.4%
+Self-instruct 67.2% 53.8% -13.4%
+Average 64.3% 50.5% -13.8%
+Table 2: Comparison of $\rho_{\text{non-trans}}$ (average preference non-transitivity) between models fine-tuned on “Raw” and “Cleaned” training sets using Qwen2.5-7B-Instruct.
+Dataset Raw Cleaned $\Delta$
+Helpful_Base 0.820 0.717 -0.103
+Vicuna 0.806 0.737 -0.069
+Oasst 0.797 0.710 -0.087
+Koala 0.804 0.718 -0.086
+Self-instruct 0.829 0.735 -0.094
+Average 0.811 0.723 -0.088
+Table 3: Comparison of $\tau_{\text{avg}}$ (average preference clarity) between models fine-tuned on “Raw” and “Cleaned” training sets using Qwen2.5-7B-Instruct.
+Human validation of discarded data quality.
+To verify the low quality of discarded data, we evaluated 100 randomly sampled instances from both “Cleaned” and “Discarded” sets across five datasets (1,000 total instances).
+Three independent annotators assessed these samples, enabling measurement of inter-annotator consistency and model-human alignment.
+Results in Table 4 indicate that human annotator consistency in the “Discard” training set averages only 34.4%, significantly lower than the 52.6% observed in the “Cleaned” training set.
+This confirms that data leading to non-transitive preferences is inherently ambiguous.
+Furthermore, model evaluation reveals that consistency between the “Discarded” training set and human majority votes averages merely 51.2%, substantially below the 80.6% achieved by the “Cleaned” training set.
+These findings provide compelling evidence that the “Discarded” training set contains low-quality preference pairs unsuitable for training models to learn consensus human preferences.
+For deeper insight into non-transitive preferences, we observed that source response pairs exhibiting non-transitivity demonstrate higher textual similarity.
+This observation suggests that preference cycles are likely to occur when the quality difference between source responses falls below the just noticeable difference (JND) (stern2010just) threshold of the evaluator LLM.
+As shown in Table 5, we quantified content similarity between response pairs by calculating Self-Bleu scores.
+Source response pairs with non-transitive preference relations consistently displayed higher Self-Bleu values than those with transitive relations, confirming that preference non-transitivity correlates with greater text similarity between comparative samples.
+HC MHA
+Dataset Cleaned Discarded Cleaned Discarded
+Help. 59% 46% 72% 45%
+Vicuna 48% 24% 78% 47%
+Oasst 56% 33% 83% 53%
+Koala 54% 41% 84% 61%
+Self. 46% 28% 86% 50%
+Average 52.6% 34.4% 80.6% 51.2%
+Table 4: Comparison of Human Evaluation Consistency and Model-Human Agreement between training sets.
+HC: Human Consistency (inter-annotator consensus); MHA: Model-Human Agreement (alignment with majority vote).
+Help.
+= Helpful_Base; Self.
+= Self-instruct.
+Pref. Helpful. Vicuna Oasst Koala Self.
+Non. 0.0903 0.1026 0.0946 0.0888 0.1135
+Trans. 0.0713 0.0918 0.0788 0.0746 0.0829
+Table 5: Comparison of source text pair Self-Bleu scores across datasets.
+Non.
+and Trans.
+indicate source pairs from non-transitive and transitive preferences; Helpful.
+= Helpful_Base; Self.
+= Self-instruct.
+Performance analysis of real-world evaluation systems.
+We compared the performance of our trained evaluator LLMs within the MT-bench evaluation framework (zheng2024judging_mt).
+We adopted the adjusted win rate calculation consistent with the MT-bench evaluation methodology.
+This metric normalizes performance by treating ties as partial wins, defined as $r_{\text{adj}}=(r_{w}+0.5\cdot r_{t})/(r_{w}+r_{l}+r_{t})$, where $r_{w}$, $r_{l}$, and $r_{t}$ represent the win rate, loss rate, and tie rate respectively.
+The adjusted rate $r_{\text{adj}}$ serves as the primary indicator for our model ranking.
+Results in Table 6 demonstrate that models fine-tuned on the “Cleaned” training sets exhibit significantly higher standard deviations in their adjusted win rates across five datasets.
+This indicates enhanced robustness of the evaluation results and better differentiation of performance differences between models.
+Dataset Help. Vicu. Oasst Koala Self.
+Cleaned 16.1% 16.6% 13.1% 14.9% 9.7%
+Raw 12.2% 13.6% 11.0% 12.3% 8.2%
+$\Delta$ +3.9% +3.0% +2.1% +2.6% +1.5%
+Table 6: Standard deviation analysis. Help. = Helpful_Base; Vicu. = Vicuna; Self. = Self-instruct.
+MHA SC
+Dataset Raw Cleaned Raw Cleaned
+Helpful_base 66.9% 66.9% 0.93 0.97
+Vicuna 65.2% 65.8% 0.97 0.98
+Oasst 66.4% 67.6% 0.93 0.93
+Koala 66.4% 66.9% 0.98 0.98
+Self-instruct 67.8% 68.3% 0.98 1.00
+Average 66.5% 67.1% 0.96 0.97
+Table 7: Comparison of Model-Human Agreement (MHA) and Spearman Correlation (SC) between fine-tuned models. Spearman Correlation quantifies the rank correlation between model and human preferences.
+Analysis of human agreement impact.
+The model’s alignment with human preferences was evaluated by calculating agreement rates and Spearman correlation coefficients on 2.5k manually annotated preference labels from AlpacaEval.
+Results in Table 7 indicate that models fine-tuned on the “Cleaned” training set consistently outperformed those trained on “Raw” training set across multiple dataset, with maximum improvements of 1.2% in human agreement and 0.04 in Spearman correlation.
+These results demonstrate that filtering ambiguous, low-quality preference data not only improves discrimination and enhances the performance of downstream evaluation systems but also increases consistency with human judgment, providing strong evidence for practical applications.
+5.2 Ablation Studies
+Effect of Different Base Models: Testing using the same steps with LLaMA3.1-8B-Instruct confirms our findings: models trained on the “Cleaned” training set consistently outperformed those trained on the “Raw” training set (Table 1).
+Data Filtering Analysis: Our “Cleaned” set retained approximately 80% of the raw data (Figure 4).
+A control experiment with random 20% filtering (Qwen-Random) produced results similar to training on the “Raw” set but with higher preference non-transitivity, confirming that our targeted approach improves data quality beyond simple reduction (Table 1).
+Prompt Format Variations: Additional experiments with prompts explicitly allowing “tie” judgments further validated our methodology.
+Detailed prompt templates and results are provided in Appendices B and C.
 6 Conclusion
-In this paper, we present a detailed analysis of preference non-transitivity in pairwise evaluations by LLM-as-a-Judge using graph theory, while measuring overall clarity through the two-dimensional structural entropy of directed graphs.
-By filtering low-quality training data that induce non-transitive preferences, our method effectively reduces preference non-transitivity, enhances overall preference clarity, and improves alignment with human judgments.
-This principled framework contributes to more reliable automated evaluation with LLMs and offers valuable insights for future research on preference consistency and robust benchmarking.
-Limitations and Future Work
-Although we have validated the effectiveness of our filtering approach on five datasets from AlpacaEval, it still cannot fully cover all real-world scenarios, and further validation across a broader range of domains is needed.
-Additionally, while our method has successfully reduced the issue of preference intransitivity, it has not been completely eliminated.
-Further efforts to minimize preference intransitivity remain an important direction for future work.
-Lastly, our study primarily focused on pairwise comparison tasks in single-turn dialogues, exploring multi-turn dialogue tasks presents an intriguing avenue for future research.
-Ethics Statement
-This research focuses on analyzing and mitigating non-transitive preferences in LLMs when used as evaluators for open-ended tasks.
-The methodology involves processing publicly available data and conducting computational experiments using existing datasets and open-source frameworks, without involving human subjects or sensitive personal information.
-Therefore, no specific ethical approval was required for this study.
-Appendix A LLM Details
-This section provides a detailed overview of the models used to construct the training and testing sets in the experiment.
-We selected a total of 21 LLMs, covering advanced closed-source models, large-scale and small-scale open-source models, as well as dense and Mixture-of-Experts (MoE) architectures.
-Specifically, the selection includes the following:
-  • OpenAI: OpenAI model series OpenAI et al. (2024).
-  • Anthropic: Claude model seriesAnthropic (2023, 2024).
-  • Mistral: Mistral model seriesJiang et al. (2024, 2023).
-  • Alibaba: Qwen model seriesQwen Team (2024a, b).
-  • Meta: LLaMA model series Meta AI (2024a, b); Touvron et al. (2023).
-  • Yi: Yi model seriesAI et al. (2025).
-  • Google: Gemini model seriesTeam et al. (2024).
-  • Other Notable Models: Vicuna 33BChiang et al. (2023), Tulu 2+DPO 70BIvison et al. (2023), and WizardLM 13BXu et al. (2024a).
-The pre-generated outputs for each model on the AlpacaEval dataset were sourced from the AlpacaEval leaderboard.
-Appendix B Additional Experimental Results
-B.1 Experimental Results of Different Prompt Forms.
-We repeated the experiment using the CoT Comparison (Tie Allowed) prompt template.
-Figure 4 illustrates the distribution of dataset sizes before and after filtering across different training sets.
-The test results are presented in Table 5.
-Across five testing sets, models fine-tuned on the "Cleaned" training set exhibit the highest preference clarity.
-In four of these testing sets (excluding Vicuna), these models also demonstrate the lowest preference non-transitivity, outperforming models fine-tuned on the "Raw" training set and the teacher model Qwen2.5-Max.
-We observe that models fine-tuned on the "Raw" training set, as well as the teacher model Qwen2.5-Max, exhibit a higher rate of ties in preferences on the Vicuna testing set.
-The prevalence of such ties—often arising from SCCs formed by cases like $(A=B)$, $(B=C)$, and $(C=A)$—significantly reduces the degree of preference non-transitivity when these SCCs are removed.
-As shown in Table 4, we compare the probability of ties among three models on the Vicuna testing set.
-Furthermore, the $\tau_{\text{avg}}^{\downarrow}$ values of models fine-tuned on the "Cleaned" training set indicate a higher overall clarity in preferences.
-Figure 4: Comparison of Data Volumes Between "Raw" and "Cleaned" Training Sets Across Different Datasets (Using the CoT Comparison (Tie Allowed) Prompt Template)
-Model tie probability
-Qwen2.5-Max 7.59%
-Qwen-Raw 10.06%
-Qwen-Cleaned 6.13%
-Table 4: Comparison of the probability of ties for models on the Vicuna testing set.
-B.2 Experimental Results of Different Base Model.
-We repeated the experiment using the LLaMA3.1-8B-Instruct model as the base model.
-The test results are presented in Table 6.
-We also compared our results with the advanced open-source Evaluator LLM Auto-J 13B, which is fine-tuned based on the LLaMA-2-13B-Chat model.
-Due to the limited context length of 2048 tokens supported by the advanced open-source evaluators PandaLM and JudgeLM, which is insufficient to cover the evaluation tasks in our testing set, we did not include them in our experiments.
-B.3 "Unseen" question validation
-We perform a detailed comparison of the performance of models fine-tuned on different training sets and the advanced Evaluator LLM on various testing sets:
-  • Results for models based on Qwen2.5-7B-Instruct are presented in Table 7.
-  • Results for models based on LLaMA3.1-8B-Instruct are presented in Table 8.
-Appendix C Evaluate Prompts
-C.1 CoT Comparison
-The CoT Comparison prompt template, identical to AlpacaEval 2.0 Li et al. (2023), is detailed in Table 9.
-C.2 CoT Comparison (Tie Allowed)
-The CoT Comparison (Tie Allowed) prompt template, as described in (Xu et al., 2025), is detailed in Table 10.
-Appendix D Under What Scenarios Are Evaluator LLMs Prone to Non-Transitive Preferences?
-Evaluator LLMs, much like human judges, tend to exhibit non-transitive preferences under certain conditions.
-First, when the quality differences between responses fall below the Evaluator LLM’s just noticeable difference (JND) threshold (a concept widely used in psychophysics to describe the minimum difference perceptible to an observer Stern and Johnson (2010)) the model may "lose evaluative focus." Analogous to how humans struggle to differentiate between highly similar colors on a gradient from yellow to orange, the Evaluator LLM may perceive adjacent responses as "almost equally good." While individual comparisons appear consistent, cumulative evaluations can yield preference cycles (e.g., $A\succ B$, $B\succ C$, $C\succ A$), reflecting underlying non-transitivity.
-Second, when responses differ significantly in stylistic dimensions, the Evaluator LLM may exhibit indecision due to the inherent conflict in multi-dimensional evaluation.
-Consider an open-ended question evaluated along three criteria: rigor, vividness, and richness:
-  • Response A is the most rigorous, moderately vivid, and least rich.
-  • Response B is the richest, moderately rigorous, and least vivid.
-  • Response C is the most vivid, moderately rich, and least rigorous.
-If overall preference is defined as outperforming another response in at least two out of three dimensions, instances such as $A\succ B$, $B\succ C$, and $C\succ A$ may arise—resembling the classic voting paradox, also known as the Condorcet paradox Fishburn (1974).
-This reflects a fundamental limitation in aggregating preferences across conflicting criteria: when each response possesses a distinct advantage, constructing a consistent global ranking becomes infeasible.
-The problem is especially pronounced in open-ended evaluations, where stylistic variance (e.g., concise and rigorous analysis vs.
-vivid and elaborate narratives) can lead to non-transitive outcomes in Evaluator LLM preferences.
-Appendix E Algorithm Detail
+We propose ELSPR, a graph-theoretic framework to analyze and mitigate non-transitivity in evaluator LLMs by modeling pairwise comparisons as a tournament graph and filtering problematic data.
+Ambiguous, low-quality preferences are a major source of non-transitive judgments, and targeted filtering improves evaluation reliability.
+Empirical results show that filtered data aligns better with human judgments and can fine-tune models to reduce non-transitive bias and structural entropy.
+This work underscores the importance of data quality and provides a scalable approach to enhance consistency on future benchmarks.
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China under Grant Nos. 62572105 and U22B2005; the Liaoning Revitalization Talents Program under Grant No. XLYC2403086.
+Appendix A Algorithm Detail
+In this section, we provide the detailed pseudocode for the filtering algorithm described in Section 3.3 of the main text.
 Algorithm 1 Filtering Strategy for Preference Data That Induces Non-Transitivity
 1:Cyclic directed graph $G=(V,E)$
-2:Cleaned dataset with transitive preference relations
-3:Decompose $G$ into strongly connected components $SCCs$: $\{SCC_{1},SCC_{2},\dots,SCC_{n}\}$
+2:Clean dataset with transitive preference relations
+3:Decompose $G$ into strongly connected components SCCsSCCs: $\{SCC_{1},SCC_{2},\dots,SCC_{n}\}$
 4:for each $SCC_{i}\in\{SCC_{1},\dots,SCC_{n}\}$ do
@@
 19:  if edge is $(v_{i}\rightarrow v_{j})$ then
-[…]
+20:   Set $\mathcal{J}(a_{i},a_{j})=\text{`lose'}$, $\mathcal{J}(a_{j},a_{i})=\text{`win'}$
 21:  else if edge is $(v_{i}\leftrightarrow v_{j})$ then
-[…]
+22:   Set $\mathcal{J}(a_{i},a_{j})=\mathcal{J}(a_{j},a_{i})=\text{`tie'}$
 23:  end if
@@
 25:Initialize empty dataset Cleaned
-26:for each data point in original dataset do
-27:     if preference relation matches $\mathcal{J}$ then
-28:         Add data point to Cleaned
-29:     end if
-30:end for
-31:return Cleaned
+26:Initialize empty dataset Discarded
+27:for each data point in original dataset do
+28:  if preference relation matches $\mathcal{J}$ then
+29:   Add data point to Cleaned
+30:  else
+31:   Add data point to Discarded
+32:  end if
+33:end for
+34:return Cleaned, Discarded
+Appendix B Additional Experimental Results
+B.1 Experimental Results of Different Prompt Forms.
+We repeated the experiment using the CoT Comparison (Tie Allowed) prompt template.
+Figure 5 illustrates the distribution of dataset sizes before and after filtering across different training sets.
+The test results are presented in Table 9.
+Across five testing sets, models fine-tuned on the “Cleaned” training set exhibit the highest preference clarity.
+In four of these testing sets (excluding Vicuna), these models also demonstrate the lowest preference non-transitivity, outperforming models fine-tuned on the “Raw” training set and the teacher model Qwen2.5-Max.
+We observe that models fine-tuned on the “Raw” training set, as well as the teacher model Qwen2.5-Max, exhibit a higher rate of ties in preferences on the Vicuna testing set.
+The prevalence of such ties—often arising from SCCs formed by cases like $(A=B)$, $(B=C)$, and (C=A)(C=A)—significantly reduces the degree of preference non-transitivity when these SCCs are removed.
+As shown in Table 8, we compare the probability of ties among three models on the Vicuna testing set.
+Furthermore, the $\tau_{\text{avg}}^{\downarrow}$ values of models fine-tuned on the “Cleaned” training set indicate a higher overall clarity in preferences.
+Figure 5: Comparison of Data Volumes Between “Raw” and “Cleaned” Training Sets Across Different Datasets (Using the CoT Comparison (Tie Allowed) Prompt Template)
+Model tie probability
+Qwen2.5-Max 7.59%
+Qwen-Raw 10.06%
+Qwen-Cleaned 6.13%
+Table 8: Comparison of tie probabilities in evaluation results produced by different evaluator LLMs on Vicuna testing set.
+B.2 “Unseen” question validation
+We perform a detailed comparison of the performance of models fine-tuned on different training sets and the advanced evaluator LLM on various testing sets:
+  • Results for models based on Qwen2.5-7B-Instruct are presented in Table 10.
+  • Results for models based on LLaMA3.1-8B-Instruct are presented in Table 11.
+Appendix C Evaluate Prompts
+C.1 CoT Comparison
+The CoT Comparison prompt template, identical to AlpacaEval 2.0, is detailed in Table 12.
+C.2 CoT Comparison (Tie Allowed)
+The CoT Comparison (Tie Allowed) prompt template, is detailed in Table 13.
+Appendix D Limitations and Future Work
+Although we have validated the effectiveness of our filtering approach on five datasets from AlpacaEval, it still cannot fully cover all real-world scenarios, and further validation across a broader range of domains is needed.
+Additionally, while our method has successfully reduced the issue of preference non-transitivity, it has not been completely eliminated.
+Further efforts to minimize preference intransitivity remain an important direction for future work.
+Lastly, our study primarily focused on pairwise comparison tasks in single-turn dialogues, exploring multi-turn dialogue tasks presents an intriguing avenue for future research.
+Appendix E Ethics Statement
+This research aims to analyze and mitigate non-transitive preferences in large language models when employed as evaluators for open-ended tasks.
+Our methodology exclusively utilizes publicly available data, existing datasets, and open-source frameworks for computational experiments.
+All human annotation revisions were conducted by graduate students under fair working conditions with appropriate compensation, ensuring both high-quality data collection and adherence to ethical standards.
+We confirm that our research procedures comply with established ethical guidelines for AI research and do not present foreseeable negative societal impacts.
 Model Helpful_Base Vicuna Oasst Koala Self-instruct
@@
 Table 9: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs ( using the CoT Comparison (Tie Allowed) prompt template).
-Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively.
-For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data.
-Model Helpful_Base Vicuna Oasst Koala Self-instruct
-$\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$ $\rho_{\text{non-trans}}^{\downarrow}$ $\tau_{\text{avg}}^{\downarrow}$
-Auto-J-13B 51.50% 0.7049 63.57% 0.7765 54.86% 0.7101 53.94% 0.7213 58.90% 0.7371
-LLaMA-Base 76.41% 0.8518 66.96% 0.8461 67.40% 0.7933 69.87% 0.8180 71.03% 0.8091
-LLaMA-Raw 59.03% 0.7654 60.18% 0.7718 58.21% 0.7595 57.22% 0.7457 59.95% 0.7830
-LLaMA-Cleaned (ours) 40.20% 0.6523 45.36% 0.6910 43.01% 0.6288 44.38% 0.6591 41.95% 0.6423
-Table 6: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs.
-LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model, while LLaMA-Raw, LLaMA-Random, and LLaMA-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively.
-For example, LLaMA-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data.
-Auto-J 13B is fine-tuned based on the LLaMA-2-13B-Chat model.
+Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw and Qwen-Cleaned denote models fine-tuned on the “Raw” and “Cleaned” training sets, respectively.
+For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the “Cleaned” training set derived from filtered Helpful_Base data.
 Model Helpful_Base Vicuna Oasst Koala Self-instruct Average
@@
 Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively.
-Specifically, Qwen-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while Qwen-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
+Specifically, Qwen-Helpful_Base-Raw is fine-tuned on the “Raw” training data generated from the Helpful_Base dataset, while Qwen-Helpful_Base-Cleaned is fine-tuned on the corresponding “Cleaned” dataset after removing non-transitive preference data.
 Model Helpful_Base Vicuna Oasst Koala Self-instruct Average
@@
 Base Model
-Auto-J-13B 51.50% 0.7049 63.57% 0.7765 54.86% 0.7101 53.94% 0.7213 58.90% 0.7371 56.55% 0.7300
 LLaMA-Base 76.41% 0.8518 66.96% 0.8461 67.40% 0.7933 69.87% 0.8180 71.03% 0.8091 70.33% 0.8237
@@
 Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively.
-Specifically, LLaMA-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while LLaMA-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
+Specifically, LLaMA-Helpful_Base-Raw is fine-tuned on the “Raw” training data generated from the Helpful_Base dataset, while LLaMA-Helpful_Base-Cleaned is fine-tuned on the corresponding “Cleaned” dataset after removing non-transitive preference data.
 System
@@
 ## Instruction
-"instruction": """{instruction}""",
+”instruction”: ”””{instruction}”””,
 ## Model Outputs
 Here are the unordered outputs from the models. Each output is associated with a specific model, identified by a unique model identifier.
-"model_identifier": "m",
-"output": """{output_1}"""
-"model_identifier": "M",
-"output": """{output_2}"""
+”model_identifier”: ”m”,
+”output”: ”””{output_1}”””
+”model_identifier”: ”M”,
+”output”: ”””{output_2}”””
 ## Task
@@
 Now is your turn.
-## Your answer: "Concise explanation" followed by "Which is best, m or M?"
+## Your answer: ”Concise explanation” followed by ”Which is best, m or M?”
 Table 12: The Chain-of-Thought Comparison prompt for pairwise comparison.
@@
 ## Instruction
-"instruction": """{instruction}""",
+”instruction”: ”””{instruction}”””,
 ## Model Outputs
 Here are the unordered outputs from the models. Each output is associated with a specific model, identified by a unique model identifier.
-"model_identifier": "m",
-"output": """{output_1}"""
-"model_identifier": "M",
-"output": """{output_2}"""
+”model_identifier”: ”m”,
+”output”: ”””{output_1}”””
+”model_identifier”: ”M”,
+”output”: ”””{output_2}”””
 ## Task
@@
 Now is your turn.
-## Your answer: "Concise explanation" followed by "Which is best, m, M or D?"
+## Your answer: ”Concise explanation” followed by ”Which is best, m, M or D?”
 Table 13: The Chain-of-Thought Comparison prompt (Tie Allowed) for pairwise comparison.